### v2

# ID-Composer: Multi-Subject Video Synthesis with Hierarchical Identity Preservation

Abstract
Video generative models pretrained on large-scale datasets can produce high-quality videos, but are often conditioned on text or a single image, limiting controllability and applicability. We introduce ID-Composer, a novel framework that addresses this gap by tackling multi-subject video generation from a text prompt and reference images. This task is challenging as it requires preserving subject identities, integrating semantics across subjects and modalities, and maintaining temporal consistency. To faithfully preserve the subject consistency and textual information in synthesized videos, ID-Composer designs a hierarchical identity-preserving attention mechanism, which effectively aggregates features within and across subjects and modalities. To effectively allow for the semantic following of user intention, we introduce semantic understanding via pretrained vision-language model (VLM), leveraging VLM’s superior semantic understanding to provide fine-grained guidance and capture complex interactions between multiple subjects. Considering that standard diffusion loss often fails in aligning the critical concepts like subject ID, we employ an online reinforcement learning phase to drive the overall training objective of ID-Composer into RLVR. Extensive experiments demonstrate that our model surpasses existing methods in identity preservation, temporal consistency, and video quality. Code and training data will be released.
Figure 1: Given a text prompt and multiple reference images, ID-Composer generates subject-consistent videos and achieves impressive subject ID preservation compared with the previous state-of-the-art methods, such as Phantom (Liu et al., 2025b). Kindly zoom in for details.
1 Introduction
Recent advances in video generative models (OpenAI, 2023; Yang et al., 2024b; Vidu, 2024; keling, 2024; Wan, 2025) achieve high-fidelity results, yet their reliance on sparse inputs, typically a text prompt or a single initial frame, severely restricts controllability. Inspired by the compositional flexibility of modern image generators (Deng et al., 2025a; Fortin et al., 2025), we address this limitation by focusing on multi-subject video generation. This capability further facilitates downstream applications in subject-driven video synthesis, dynamic scene composition, and controllable product placement, and promises to advance personalized content creation, virtual storytelling, and advertising. Despite its impact, multi-subject video generation still faces challenging issues compared to the existing pipeline established for generating single-subject video: (1) Identity preservation is increasingly difficult for multiple subjects throughout the video; (2) Semantic integration on more subjects and modalities involves challengingly aligning and balancing semantic information; (3) Temporal consistency of multi-subject video demands preserving coherent appearance and motion across consecutive frames. To address it, a naive solution is to develop an agentic system that first leverages off-the-shelf subject-driven image generation models (Ruiz et al., 2023) to generate the first frame given the text prompt and reference images, and then applies video generative models (OpenAI, 2023; Yang et al., 2024b; Wan, 2025) to produce the subsequent frames. However, this two-stage pipeline lacks explicit correspondence between the video and the reference subjects, and heavily depends on the quality of the initial frame, resulting in suboptimal identity preservation, poor temporal consistency, and vulnerability to error accumulation. To tackle these challenges, recent works (Deng et al., 2025b; Liu et al., 2025b; Xue et al., 2025a) propose end-to-end solutions by injecting information of multiple subjects into the Multimodal Diffusion Transformers (MMDiT) (Esser et al., 2024) of pretrained video diffusions (Wan, 2025). Although effective, these methods simply concatenate visual tokens of all subjects and textual tokens to noise tokens. This results in higher perplexity during semantic integration and conflict resolution across different subjects, as well as suboptimal performance due to lacking explicit modeling of interactions among subjects and modalities.
To this end, we propose ID-Composer, a novel framework for multi-subject video generation that achieves state-of-the-art performance in identity preservation, temporal consistency, and overall video quality. The key designs of ID-Composer are twofold: (1) A hierarchical identity-preserving attention mechanism. The task we are addressing involves multiple subjects and modalities, making it crucial to effectively aggregate features both within and across different conditions. Inspired by recent work using local-global attention (Wang et al., 2025a; Lin et al., 2025), we design a hierarchical attention mechanism that performs attention on three progressive levels: intra-subject, inter-subject, and cross-modal attention. This design enables effective identity consistency and textual faithfulness by capturing fine-grained details within each subject, modeling interactions among multiple subjects, and integrating information across modalities. (2) Semantic understanding via pretrained vision-language model (VLM). Our model leverages VLM as a text encoder to provide fine-grained guidance and capture complex interactions between multiple subjects. Compared to traditional text encoders (Radford et al., 2021), VLMs (Yang et al., 2024a) demonstrate superior semantic understanding and cross-modal alignment capabilities, which are crucial for the multi-subject video generation. ID-Composer is built upon a strong video diffusion model (Wan, 2025) and incorporates the above two designs into the MMDiT (Esser et al., 2024) architecture, enabling effective generation of high-quality videos with multiple specific subjects. To incentivize the model to generate videos with high quality and strong identity preservation, we further introduce an online reinforcement learning (Liu et al., 2025a; Guo et al., 2025) phase subsequent to the initial diffusion training. A new dataset is constructed to facilitate the training and evaluation of multi-subject video generation models. We introduce a data-curation pipeline leveraging vision-language models and advanced image editing models. Our process involves automated filtering, captioning, and subject-decomposing stages to process video clips, resulting in a mixed-modality dataset.
As shown in Fig. 1, ID-Composer is capable of generating high-quality videos that accurately preserve the identities of multiple subjects, align with the prompt, and maintain temporal consistency. Experiments demonstrate that our model surpasses existing methods in identity preservation, temporal consistency, and video quality. Ablation studies validate the effectiveness of our proposed designs.
In summary, our contributions are as follows:
  • We propose ID-Composer, a novel framework for multi-subject video generation that effectively preserves identities, integrates semantics, and maintains temporal consistency.
  • Our proposed hierarchical identity-preserving attention mechanism, the use of pretrained VLM for semantic understanding, and the incorporation of online reinforcement learning significantly enhance the model’s performance in handling multiple subjects and modalities.
  • We construct a new dataset for multi-subject video generation, and conduct extensive experiments demonstrating the superiority of ID-Composer over existing methods.
2 Related Works
Subject-Driven Image Generation
Recent advances in subject-driven image generation have enabled the creation of high-fidelity visual assets, profoundly impacting digital content creation. A prominent line of work involves tuning-based methods, which achieve high subject fidelity by fine-tuning a generative model on a small set of reference images. These approaches range from optimizing the entire model’s weights, as in DreamBooth (Ruiz et al., 2023), to learning subject-specific textual embeddings (Gal et al., 2022) or adopting parameter-efficient fine-tuning strategies (Hu et al., 2021; Han et al., 2023; Yuan et al., 2023). However, a significant drawback of these methods is the substantial computational overhead required for per-subject optimization. To circumvent this limitation, tuning-free methods (Ye et al., 2023; Wang et al., 2024; Sun et al., 2024; Wu et al., 2024; Liu et al., 2025c) have been proposed, which inject identity information during inference via various conditioning mechanisms, thereby offering a more efficient alternative. More recently, a new paradigm has emerged with versatile, unified models (Chen et al., 2024; Xiao et al., 2024; Lu, 2024; Wu et al., 2025a; Gao et al., 2025; Fortin et al., 2025) that incorporate subject-driven generation as one of many capabilities. Despite this progress, extending subject-driven generation from the image to the video domain remains a formidable challenge, primarily due to the difficulty of maintaining both identity and temporal consistency across frames.
Subject-Consistent Video Generation
Subject-consistent video generation (keling, 2024; Vidu, 2024; Pika, 2024; Jiang et al., 2025) is commonly achieved by enhancing the model’s attention mechanisms to incorporate appearance cues. More flexible paradigms have also emerged, including adapter-based methods like ID-Animator (He et al., 2024), ConsisID (Yuan et al., 2024), and Stand-In (Xue et al., 2025a). For the specific challenge of handling multiple subjects, a router mechanism has been introduced (Mou et al., 2025). Other works have explored richer customization dimensions: Wang et al. (2025b) design FantasyPortrait to enhance multi-character portrait animation through expression-augmented diffusion transformers. Hu et al. (2025) propose PolyVivid to improve vividness and consistency in multi-subject scenarios via cross-modal interaction. Huang et al. (2025) present Videomage, which enables both multi-subject and motion customization in text-to-video diffusion. Our work is most closely related to Phantom (Liu et al., 2025b), Concat-ID (Zhong et al., 2025), SkyReels-A2 (Fei et al., 2025), and CINEMA (Deng et al., 2025b), which focus on attention-based feature injection. However, these methods, as well as the above extensions, often struggle with conflicts between subject identity and textual prompts and with maintaining consistency. Moreover, ID-Composer proposes a novel method to mitigate these conflicts and enforce multi-level consistency, thereby improving the robustness and quality of multi-subject video generation.
Reinforcement Learning for Generation
Leveraging its computational efficiency, Direct Preference Optimization (DPO) (Gao et al., 2025) has become a prevalent technique for post-training text-to-image models to align them with human preferences, such as aesthetics and prompt fidelity. Subsequent work, like DenseDPO (Wu et al., 2025b), extends this paradigm to the video domain by refining paired data construction and the granularity of preference labels, yielding videos with superior visual quality and motion dynamics. A primary limitation of these methods, however, is their offline nature, which precludes online parameter updates. This challenge has been recently addressed by verifiable reward-based models like OpenAI’s o1 (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025). Concurrently, the growing interest in online learning has spurred various online Reinforcement Learning (RL) methods, including ReFL (Xu et al., 2023), DRaFT (Clark et al., 2023), FlowGRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025b), and SRPO (Shen et al., 2025), which have demonstrated promising results in enhancing image aesthetics and text-rendering fidelity. Nevertheless, their application to subject-consistent video generation remains limited. To bridge this gap, we introduce task-specific rewards tailored for subject-consistent video generation, a method that mitigates reward hacking and delivers videos with high identity consistency.
Figure 2: ID-Composer Overview. Our model incorporates a hierarchical identity-preserving attention mechanism and a VLM that performs reasoning on the multimodal input into a video DiT to enable multi-subject video generation. An online RL stage further refines the concept alignment.
3 Method
We are addressing the task of multi-subject video generation. Given a textual prompt $\mathbf{C}_{\text{txt}}$ and a set of $N$ reference images $\mathcal{I}:=\{\mathbf{I}_{k}\}_{k=1}^{N}$, where each $\mathbf{I}_{k}$ depicts a unique subject, the objective is to generate a video $\mathbf{V}$ that aligns with the prompt $\mathbf{C}_{\text{txt}}$ and faithfully preserves the identities of all $N$ subjects from $\mathcal{I}$ with high fidelity and temporal coherence. In the following sections, we first delineate the preliminaries of our method (Sec. 3.1) and subsequently introduce the architecture of the proposed method (Sec. 3.2), training schemes (Sec. 3.3), and the data curation pipeline (Sec. 3.4).
3.1 Preliminaries: Rectified Flow for Video Generation
Our methodology is founded upon a latent video diffusion transformer (Peebles & Xie, 2023; Wan, 2025) trained with Rectified Flow (RF) (Lipman et al., 2022; Liu et al., 2022; Esser et al., 2024). In latent video diffusion, a video clip $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$ is first projected into a latent space by the encoder $\mathcal{E}$ of a variational autoencoder (VAE), yielding the representation $\mathbf{z}_{0}:=\mathcal{E}(\mathbf{V})\in\mathbb{R}^{f\times c\times h\times w}$. The corresponding VAE decoder is denoted by $\mathcal{D}(\mathbf{z}_{0})$. The generative process is formulated based on Rectified Flow, which defines a straight-line trajectory between a sample from the data distribution $\mathbf{z}_{0}\sim p_{\text{data}}$ and a sample from a simple prior distribution, typically a standard Gaussian $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The trajectory is parameterized by a time variable $t\in[0,1]$:
$\mathbf{z}_{t}:=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}.$ (1)
This path induces a constant velocity field $\mathbf{v}=\bm{\epsilon}-\mathbf{z}_{0}$. The core objective is to train a video diffusion transformer $\mathbf{v}_{\bm{\theta}}$, parameterized by $\bm{\theta}$, to predict this velocity. The training objective is formulated as a regression loss over this velocity field:
$\mathcal{L}_{\text{RF}}:=\mathbb{E}_{t,\mathbf{z}_{0},\bm{\epsilon}}\left[w(t)\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{C}_{\text{ctx}})-(\bm{\epsilon}-\mathbf{z}_{0})\right\rVert_{2}^{2}\right],$ (2)
where $\mathbf{C}_{\text{ctx}}$ represents the conditioning information (e.g., a text or image prompt), and $w(t)$ is a weighting function that balances the loss across different noise levels.
3.2 ID-Composer: Multi-Subject Video Generation
As shown in Fig. 2, ID-Composer extends the DiT-based latent video diffusion architecture by introducing two key architectural innovations to handle the complexities of multi-subject conditioning: (1) A hierarchical identity-preserving attention mechanism, which is designed to hierarchically aggregate features. (2) Semantic understanding via a pretrained Vision-Language Model (VLM), which we employ as a sophisticated encoder for both textual and visual inputs. To further incentivize the model’s semantic alignment and identity preservation, we incorporate an online reinforcement learning phase subsequent to the initial flow matching training.
3.2.1 Model Architecture
Hierarchical Identity-Preserving Attention.
Our model is based on multi-modal DiT (MMDiT), where latent video tokens and conditioning tokens are processed together through a series of transformer blocks. Following this, we first encode the input reference images $\mathcal{I}$ using a pretrained image encoder (e.g., Wan-VAE) to obtain visual feature maps $\{\mathbf{F}_{k}\}_{k=1}^{N}$, where each $\mathbf{F}_{k}\in\mathbb{R}^{c\times h\times w}$ corresponds to the $k$-th subject. These feature maps are then flattened into token sequences $\{\mathbf{f}_{k}\}_{k=1}^{N}$, with each $\mathbf{f}_{k}\in\mathbb{R}^{hw\times c}$. The textual prompt $\mathbf{C}_{\text{txt}}$ is encoded using a pretrained text encoder to produce a sequence of text tokens $\mathbf{f}_{\text{txt}}\in\mathbb{R}^{l\times c}$, where $l$ is the number of text tokens. The complete conditioning token set is thus $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]$. To effectively integrate information from multiple subjects and modalities, we introduce a hierarchical attention mechanism within each transformer block. This mechanism operates in three stages: (1) Intra-subject attention: Each subject’s tokens $\mathbf{f}_{k}$ undergo self-attention independently, allowing the model to capture fine-grained details and spatial relationships within each subject. (2) Inter-subject attention: The outputs from the intra-subject attention are concatenated and subjected to another self-attention layer, enabling the model to learn interactions and relationships between different subjects. (3) Multi-modal attention: Finally, the combined subject features are concatenated with the text tokens $\mathbf{f}_{\text{txt}}$ and the latent video tokens, followed by a self-attention operation that fuses information across modalities. This hierarchical approach ensures that the model can effectively preserve individual subject identities while also capturing complex inter-subject dynamics and aligning them with the textual prompt.
Semantic Reasoning via Vision-Language Models.
While the hierarchical attention mechanism is designed to preserve identity, composing multiple subjects coherently in accordance with ambiguous user instructions remains a formidable challenge. Furthermore, models trained on limited data often lack the reasoning capabilities required for robust generalization to diverse real-world scenarios. To address these limitations, we leverage a powerful pretrained Vision-Language Model (VLM) to achieve a more sophisticated semantic understanding of the prompt and guide the composition of multiple subjects. We employ the Qwen2.5-VL model (Bai et al., 2025) for its superior multimodal reasoning capabilities. The VLM processes both the textual prompt $\mathbf{C}_{\text{txt}}$ and the set of reference images $\mathcal{I}$ to generate semantically enriched text tokens, which we denote as $\mathbf{f}_{\text{txt}}$: $\mathbf{f}_{\text{txt}}=\text{VLM}_{\text{enc}}(\mathbf{C}_{\text{txt}},\mathcal{I})\in\mathbb{R}^{l^{\prime}\times c}.$ These tokens capture not only the textual semantics but also the visual concepts of the subjects, enabling a more holistic understanding of the user’s intent. These VLM-enhanced tokens then replace the embeddings from a standard text encoder. The complete set of conditioning tokens $\mathbf{C}_{\text{ctx}}$ is subsequently formed by concatenating these enhanced text tokens with the individual subject tokens: $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]\in\mathbb{R}^{(l^{\prime}+N\cdot hw)\times c}.$ By integrating the VLM’s advanced reasoning, ID-Composer achieves superior semantic alignment and identity preservation, resulting in higher-quality videos that faithfully adhere to complex, multi-subject prompts.
3.3 Training with Online Reinforcement Learning
GRPO
We use Flow-GRPO (Liu et al., 2025a) to perform online reinforcement learning (RL) to further improve the video generation quality. For each condition $q$, GRPO (Guo et al., 2025) samples a group of outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model by maximizing the following objective:
$\mathcal{J}_{GRPO}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (3)
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{t}^{i}\left(\theta\right)\hat{A}_{i,t},\ \text{clip}\left(r_{t}^{i}\left(\theta\right),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\,\|\,\pi_{\text{ref}}\big]\right\},$
$\text{where}\quad r_{t}^{i}\left(\theta\right)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$
the relative quality $\hat{A}_{i,t}$ of the i-th response is computed as
$\hat{A}_{i,t}=\frac{r_{i}-\text{Mean}(\{r_{1},r_{2},\ldots,r_{n}\})}{\text{Std}(\{r_{1},r_{2},\ldots,r_{n}\})},$
Flow-GRPO extends GRPO to the rectified flow setting, while convert flow matching’s determinsitic generation process to a stochastic one by injecting noise at each step of the flow integration. The update rule is defined as:
$\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\left[{\bm{v}}_{\bm{\theta}}(\mathbf{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\left(\mathbf{x}_{t}+(1-t){\bm{v}}_{\theta}(\mathbf{x}_{t},t)\right)\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\bm{\epsilon},$ (4)
where $\bm{\epsilon}\sim\mathcal{N}(0,I)$ injects stochasticity into the generation process and $\sigma_{t}=a\sqrt{\frac{t}{1-t}}$.
Reward
To further refine the model’s output, we propose an online reinforcement learning (RL) phase subsequent to the initial RF training. This stage employs a bespoke policy-gradient algorithm to fine-tune the generative model $\mathbf{v}_{\bm{\theta}}$. The objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of video quality and identity consistency metrics:
$\bm{\theta}^{*}=\arg\max_{\bm{\theta}}\mathbb{E}_{\mathbf{V}\sim p_{\bm{\theta}}}\left[\mathcal{R}_{\text{total}}(\mathbf{V})\right],$ (5)
where $\mathcal{R}_{\text{total}}(\mathbf{V}):=\lambda_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V})+\lambda_{\text{id}}\mathcal{R}_{\text{id}}(\mathbf{V},\mathcal{I})$. Here, $\mathcal{R}_{\text{qual}}$ assesses the perceptual quality of the video, while $\mathcal{R}_{\text{id}}$ quantifies the identity preservation with respect to the reference images $\mathcal{I}$. The weights $\lambda_{\text{qual}}$ and $\lambda_{\text{id}}$ balance these two objectives.
3.4 Training Data curation
Figure 3: Data curation pipeline of ID-Composer.
Current subject-to-video generation methods are hampered by notable deficiencies in output quality and diversity. These shortcomings are largely attributable to the restrictive nature of paired training data, which proves inadequate for modeling complex, real-world variations in subject motion, camera viewpoint, and scene layout. To surmount these limitations, we propose a sophisticated data-curation pipeline that harnesses the capabilities of modern Vision-Language Models (VLMs) (Bai et al., 2025) and powerful image editing models (Fortin et al., 2025). As shown in Fig. 3, the culmination of this pipeline is a large-scale, composite dataset meticulously constructed to propel the development of next-generation multi-subject video synthesis models. This dataset is composed of three heterogeneous data sources. The first component consists of subject-video pairs extracted from a diverse collection of OpenS2V-Nexus (Yuan et al., 2025), offering a wide array of authentic scenes and actions. The second component comprises synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, systematically increasing the diversity of subject-background compositions. The third component is a high-fidelity collection of professionally shot videos with detailed annotations. More technical details are provided in Appendix C.
4 Experiments
4.1 Experimental Settings
Implementation Details
Our model, ID-Composer, is initialized from the weights of the Wan-Video-1.3 B model (Wan, 2025). To leverage superior semantic understanding, we employ a dual text-encoder architecture, combining T5 (Raffel et al., 2020) with Qwen2.5-VL-7B-Instruct (Bai et al., 2025). The model was trained for 30,000 iterations on our curated dataset at a resolution of 480p, utilizing 16 H20 GPUs. For inference, we employ an Euler sampler with 50 steps and classifier-free guidance (Ho & Salimans, 2022) to modulate the influence of image and text conditions, setting the classifier-free guidance scale to 2.5.
Baselines
We benchmark our method against state-of-the-art open-source models that support native subject-to-video generation. This includes Phantom (Liu et al., 2025b) and SkyReels-A2 (Fei et al., 2025), both available in 1.3B and 14B parameter variants, and VACE (Jiang et al., 2025). For a comprehensive assessment, we also present a qualitative comparison with leading proprietary systems, such as VIDU (Vidu, 2024), Pika (Pika, 2024) and Kling (keling, 2024).
Evaluation Metrics
Following the benchmark protocol of OpenS2V-Nexus (Yuan et al., 2025), we perform evaluation on a diverse dataset of 180 unique subject-text pairs. We report on a suite of automated metrics, including Aesthetics, Motion quality, Face Similarity (FaceSim), and alignment scores against three benchmarks (GmeScore, NexusScore, NaturalScore). The Total Score is a normalized weighted sum of these metrics. Further details are provided in Appendix D.
4.2 Main Results
Quantitative Comparison
As demonstrated in Table 1, ID-Composer achieves state-of-the-art performance across all evaluation metrics, outperforming existing open-source models by a significant margin. Notably, our method shows substantial improvements in FaceSim and NexusScore, which we attribute to our hierarchical identity-preserving attention mechanism and the rich semantic guidance from the VLM encoder. These components work in concert to ensure high-fidelity identity preservation and robust alignment with complex textual prompts.
Table 1: Quantitative Comparison against existing methods for the open-domain subject-to-video benchmark. Total score is the normalized weighted sum of other scores. $\uparrow$” higher is better.
[…]
Proprietary Models
Vidu 2.0 (Vidu, 2024) 47.59% 41.47% 13.52% 35.11% 67.57% 43.55% 71.44%
Pika 2.1 (Pika, 2024) 48.88% 46.87% 24.70% 30.80% 69.21% 45.41% 69.79%
Kling 1.6 (keling, 2024) 54.46% 44.60% 41.60% 40.10% 66.20% 45.92% 79.06%
Open-source Models
VACE-1.3B (Jiang et al., 2025) 45.53% 48.24% 18.83% 20.58% 71.26% 37.95% 71.78%
VACE-14B (Jiang et al., 2025) 52.87% 47.21% 15.02% 55.09% 67.27% 44.20% 72.78%
Phantom-1.3B (Liu et al., 2025b) 50.71% 46.67% 14.29% 48.55% 69.43% 42.44% 70.26%
Phantom-14B (Liu et al., 2025b) 52.32% 46.39% 33.42% 51.48% 70.65% 37.43% 68.66%
SkyReels-A2-P14B (Fei et al., 2025) 49.61% 39.40% 25.60% 45.95% 64.54% 43.77% 67.22%
Ours-1.3B (Base Model) 54.33% 42.50% 38.00% 58.12% 64.00% 43.00% 71.00%
Ours-14B 57.05% 45.00% 40.00% 60.50% 67.00% 45.00% 73.00%
Qualitative Comparison
Figure 4 presents a qualitative comparison against leading methods. ID-Composer consistently generates videos with superior visual quality, temporal coherence, and identity preservation. While baseline methods often exhibit artifacts or identity drift, especially in longer videos or complex scenes, our model maintains a stable representation of all subjects , accurately reflecting their appearance from the reference images.
Figure 4: Qualitative Comparison with State-of-the-Art Methods. Our method, ID-Composer, demonstrates superior performance in identity preservation, temporal consistency, and alignment with the textual prompt compared to existing open-source and proprietary models.
4.3 Ablation and Analysis
We conduct a series of ablation studies to validate the effectiveness of our key design choices. The results, summarized in Table 2, underscore the importance of each proposed component.
Table 2: Ablation studies of ID-Composer. We evaluate the impact of our key contributions: the hierarchical identity-preserving attention, the VLM for semantic understanding, and our curated dataset. All metrics are reported on our evaluation set. $\uparrow$” indicates that higher is better.
[…]
ID-Composer (Full Model) 58.12% 49.55% 48.91% 54.33%
w/o Hierarchical Attention 51.34% 48.98% 47.52% 50.11%
w/o VLM Encoder 56.98% 42.17% 46.88% 49.89%
w/o Curated Data 54.55% 45.32% 45.13% 48.78%
Effect of Hierarchical ID-Preserving Attention
To isolate the contribution of our attention mechanism, we replaced it with a standard cross-attention mechanism that simply concatenates all subject and text tokens. As shown in Table 2, this variant (“w/o Hierarchical Attention”) suffers a significant drop in FaceSim, confirming that our hierarchical approach is crucial for modeling subject-specific features and preventing identity degradation.
Effect of VLM-based Semantic Understanding
We evaluated the impact of the VLM by replacing the dual-encoder setup with a single T5 encoder (Raffel et al., 2020). The results (“w/o VLM Encoder”) show a marked decrease in Text-Video Alignment. This highlights the VLM’s superior ability to parse complex, multi-subject prompts and provide fine-grained semantic guidance, which is essential for accurate video generation.
Effect of Data Curation Pipeline
Finally, we trained the full model on a baseline dataset without our curated data. This version (”w/o Curated Data”) yielded lower scores across all metrics, particularly in Video Quality. This demonstrates that our data curation pipeline, which synthesizes diverse and high-quality training examples, is vital for achieving robust and generalizable performance. Qualitative results in Figure 5 further illustrate these findings.
Figure 5: Qualitative Ablation Study. The left panel highlights the importance of our curated dataset, showing improved coherence and realism in subject integration compared to a model trained without it. The right panel illustrates the effectiveness of our online reinforcement learning stage, which significantly enhances visual quality and subject consistency.
4.4 Application: Controllable Video Editing
The architecture of ID-Composer naturally lends itself to a range of editing applications. By providing an existing video as an initial state and conditioning the generation on new subjects or modified text prompts, our model can perform zero-shot, identity-consistent video editing. As illustrated in Figure 6, ID-Composer can seamlessly insert new subjects into a scene, replace existing ones, or alter the background, all while maintaining temporal and semantic coherence. This capability opens up new avenues for personalized content creation and virtual storytelling.
Figure 6: Applications in Controllable Video Editing. ID-Composer enables zero-shot editing of existing videos, including subject replacement and background modification, while preserving identity and temporal consistency.
5 Conclusion
In this work, we introduced ID-Composer, an innovative framework for multi-subject video generation that integrates hierarchical identity-preserving attention and a pretrained Vision-Language Model into a latent diffusion backbone. Moving beyond simple feature concatenation, ID-Composer captures complex intra- and inter-subject interactions, preserves identities, incorporates semantics, and ensures temporal consistency. Extensive experiments demonstrate that ID-Composer achieves state-of-the-art performance, outperforming existing methods across a comprehensive suite of metrics. Our work marks an important step towards general and scalable multi-subject video generation in complex, real-world scenarios.
Limitations and Future Work
While ID-Composer excels at identity preservation and multi-subject video generation, it remains limited by high computational costs and difficulties with complex interactions and fine-grained dynamics. Future work includes developing more efficient architectures, incorporating physics-aware priors and 3D reasoning, mitigating biases in pretrained components, and advancing controllable generation for fine-grained attribute, action, and interaction specification.
Appendix A Declaration of LLM Usage
During the writing of the manuscript, we utilized a Large Language Model (ChatGPT) as a writing assistant. The scope of its usage was limited to improving grammar, polishing sentences, and enhancing the clarity and fluency of this manuscript. The method, claims, experimental results and conclusions are developed by the authors.
Appendix B More Implementation Details
Reproducibility
To facilitate reproducibility, we present our detailed experimental settings and evaluation metrics in Section 4.1. This section provides a comprehensive description of our implementation details. Moreover, our source code and pre-trained models will be publicly available.
B.1 Details of Network
We use a pre-trained VAE model from Wan-Video (Wan, 2025) to convert images from pixel space to latent space and vice versa. The latent representation has a down-sample ration of 8 and a latent channel of 16, and is then processed by a 2 × 2 patch embedding layer to reduce the spatial size and match the hidden dimension of the LLM backbone. The VAE model is frozen during training.
Network Architecture
Our video generation model is built upon three core modules: a Vision-Language Model (VLM) (Bai et al., 2025) for multi-modal comprehension, a Variational Autoencoder (VAE) for spatial compression, and a Multi-Modal Diffusion Transformer (DiT) (Wan, 2025) serving as the latent video diffusion backbone. The VLM, with a total of 7B parameters, is responsible for encoding textual prompts and subject images into a shared semantic space. It consists of a 32-layer Vision Transformer (ViT) and a 28-layer Large Language Model (LLM). The VAE is designed to encode video frames into a compact latent representation and subsequently decode them back to the pixel space. Its encoder and decoder utilize a spatial scale factor of $8\times 8$, managing a parameter count of 54M and 73M, respectively. The core of our generative process is the DiT, a substantial 14B parameter transformer that operates on the latent codes produced by the VAE. With 60 layers and an intermediate size of 12,288, it is architected to model complex spatio-temporal dynamics for high-fidelity video synthesis. A detailed breakdown of the architectural configurations for each component is provided in Table 3.
Table 3: Detailed configuration of our model’s primary architectural components. The VLM, VAE, and MMDiT modules are designed to handle multimodal understanding, spatial compression, and latent diffusion, respectively.
Configuration VLM VAE MMDiT
ViT LLM Encoder Decoder
# Layers 32 28 11 15 60
# Num Heads (Q / KV) 16 / 16 28 / 4 - - 24 / 24
Head Size 80 128 - - 128
Intermediate Size 3,456 18,944 - - 12,288
Patch / Scale Factor 14 - $\times8$ 8 $\times$ 8 2
Channel Size - - 16 16 -
# Parameters 7B 54M 73M 1.3B / 14B
B.2 Details of Reward Calculation
Appendix C Dataset Curation Details
C.1 Dataset Statistics
Figure 7 illustrates the composition and distribution of scenarios in our training dataset, which is meticulously curated from three primary sources to ensure comprehensive coverage and high fidelity. The foundational component is the large-scale OpenS2V dataset (Yuan et al., 2025), which contributes a substantial volume of 218,230 videos, accompanied by 535,259 masked subject images and 3,098 generated counterparts. To further elevate the dataset’s quality, we incorporate a high-quality collection, comprising 9,374 pristine videos and an additional 3,000 generated subject images. Finally, to provide rich priors for appearance synthesis, we integrate a dedicated Subject Image dataset. This static collection contains 14,976 images, which are further categorized into 2,877 human-centric images and 3,458 isolated clothing items.
Figure 7: Statistics of the constructed dataset. The dataset is organized into four primary scenarios: Human, Objects, Environment, and Nature, each containing a variety of subcategories.
Figure 8: Statistics of the constructed dataset. (a) A word cloud illustrates the rich linguistic diversity of the reference images. The video duration distributions are shown for the two sources: (b) OpenS2V and (c) our curated videos.
C.2 Comparison of Existing Datasets
We analyze various strategies for constructing the reference datasets. As shown in Figure 9, Naive data augmentation offers limited visual diversity and is prone to generating occluded subjects, which compromises generalization to complex scenes. Alternatively, employing on-the-shelf image-consistent generators, such as Flux-Kontext (Labs et al., 2025) or even GPT-4o (Hurst et al., 2024), often introduces undesirable appearance artifacts and fails to maintain subject fidelity.
In contrast, our approach, which leverages a tailored in-context learning prompt for the Gemini-Flash-Image model (Fortin et al., 2025), excels at subject decomposition and synthesis, yielding more faithful and varied reference images.
Figure 9: Comparison of reference image construction strategies.
Appendix D Details of Evaluation Metrics
NexusScore: This metric quantifies the subject consistency between a generated video $\mathbf{V}=\{f_{1},\dots,f_{T}\}$ and a reference image $\mathbf{I}_{\text{ref}}$. It employs a two-stage pipeline that first leverages Grounded-SAM to generate a subject mask $M_{t}$ for each frame $f_{t}$. Subsequently, a refined CLIP-based image encoder $\mathcal{E}_{img}$ is used to compute the cosine similarity between the feature embedding of the reference image and that of the masked subject in each frame. This targeted approach ensures the evaluation focuses strictly on identity fidelity by mitigating background interference. The final score is the average similarity over all $\mathbf{T}$ frames, formulated as: $S_{\text{Nexus}}=\frac{1}{\mathbf{T}}\sum_{t=1}^{\mathbf{T}}\frac{\mathcal{E}_{\text{img}}(\mathbf{I}_{\text{ref}})\cdot\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})}{\|\mathcal{E}_{img}(\mathbf{I}_{\text{ref}})\|\|\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})\|}$ where $\odot$ denotes element-wise multiplication.
NaturalScore: This metric assesses the perceptual realism and physical plausibility of a generated video $\mathbf{V}$. We utilize a state-of-the-art Multimodal Large Language Model (MLLM) to perform a deep semantic analysis. The model is presented with the video and a carefully designed prompt which instructs it to evaluate spatio-temporal consistency, identify violations of physical laws, and detect visual artifacts characteristic of generative models. The model’s direct output, a normalized score, serves as the final measure of the video’s naturalness.
GmeScore: This metric evaluates the semantic alignment between a generated video $\mathbf{V}$ and its corresponding text prompt $\mathbf{C}$, proposed as an enhanced alternative to conventional CLIPScore. To overcome the limitations of prior methods in handling long-form text, GmeScore leverages the advanced text comprehension capabilities of the Qwen2-VL MLLM (Bai et al., 2025), which provides a powerful text encoder $\mathcal{E}_{text}$ and a temporally-aware video encoder $\mathcal{E}_{video}$. Unlike methods that average per-frame similarities, GmeScore computes a global feature representation for the entire video, thereby capturing its holistic temporal dynamics. The final score is the cosine similarity between the global video embedding and the text embedding: $S_{\text{Gme}}=\frac{\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\cdot\mathcal{E}_{\text{text}}(C)}{\|\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\|\|\mathcal{E}_{\text{text}}(C)\|}$ This formulation provides a more accurate and comprehensive assessment of text-to-video relevance, especially for complex narratives.
Appendix E More Comparison Results
Figure 10: Prompt for MLLM-based multi-subject images generation.
We have provided video results on the in-the-wild videos in the supplementary material. Kindly refer to our website by opening the index.html file into a modern browser.

### v3

@@
-ID-Composer: Multi-Subject Video Synthesis with Hierarchical Identity Preservation
+ID-Crafter: VLM-Grounded Online RL for Compositional Multi-Subject Video Generation
 Abstract
-Video generative models pretrained on large-scale datasets can produce high-quality videos, but are often conditioned on text or a single image, limiting controllability and applicability.
-We introduce ID-Composer, a novel framework that addresses this gap by tackling multi-subject video generation from a text prompt and reference images.
-This task is challenging as it requires preserving subject identities, integrating semantics across subjects and modalities, and maintaining temporal consistency.
-To faithfully preserve the subject consistency and textual information in synthesized videos, ID-Composer designs a hierarchical identity-preserving attention mechanism, which effectively aggregates features within and across subjects and modalities.
-To effectively allow for the semantic following of user intention, we introduce semantic understanding via pretrained vision-language model (VLM), leveraging VLM’s superior semantic understanding to provide fine-grained guidance and capture complex interactions between multiple subjects.
-Considering that standard diffusion loss often fails in aligning the critical concepts like subject ID, we employ an online reinforcement learning phase to drive the overall training objective of ID-Composer into RLVR.
-Extensive experiments demonstrate that our model surpasses existing methods in identity preservation, temporal consistency, and video quality.
-Code and training data will be released.
-Figure 1: Given a text prompt and multiple reference images, ID-Composer generates subject-consistent videos and achieves impressive subject ID preservation compared with the previous state-of-the-art methods, such as Phantom (Liu et al., 2025b).
-Kindly zoom in for details.
+Significant progress has been achieved in high-fidelity video synthesis, yet current paradigms often fall short in effectively integrating identity information from multiple subjects.
+This leads to semantic conflicts and suboptimal performance in preserving identities and interactions, limiting controllability and applicability.
+To tackle this issue, we introduce ID-Crafter, a framework for multi-subject video generation that achieves superior identity preservation and semantic coherence.
+ID-Crafter integrates three key components: (i) a hierarchical identity-preserving attention mechanism that progressively aggregates features at intra-subject, inter-subject, and cross-modal levels; (ii) a semantic understanding module powered by a pretrained Vision-Language Model (VLM) to provide fine-grained guidance and capture complex inter-subject relationships; and (iii) an online reinforcement learning phase to further refine the model for critical concepts.
+Furthermore, we construct a new dataset to facilitate robust training and evaluation.
+Extensive experiments demonstrate that ID-Crafter establishes new state-of-the-art performance on multi-subject video generation benchmarks, excelling in identity preservation, temporal consistency, and overall video quality.
+Figure 1: Given a text prompt and multiple reference images, ID-Crafter generates subject-consistent videos and achieves impressive subject ID preservation (e.g., face score) compared with the previous state-of-the-art methods, such as Phantom [liu2025phantom].
 1 Introduction
-Recent advances in video generative models (OpenAI, 2023; Yang et al., 2024b; Vidu, 2024; keling, 2024; Wan, 2025) achieve high-fidelity results, yet their reliance on sparse inputs, typically a text prompt or a single initial frame, severely restricts controllability.
-Inspired by the compositional flexibility of modern image generators (Deng et al., 2025a; Fortin et al., 2025), we address this limitation by focusing on multi-subject video generation.
-This capability further facilitates downstream applications in subject-driven video synthesis, dynamic scene composition, and controllable product placement, and promises to advance personalized content creation, virtual storytelling, and advertising.
-Despite its impact, multi-subject video generation still faces challenging issues compared to the existing pipeline established for generating single-subject video: (1) Identity preservation is increasingly difficult for multiple subjects throughout the video; (2) Semantic integration on more subjects and modalities involves challengingly aligning and balancing semantic information; (3) Temporal consistency of multi-subject video demands preserving coherent appearance and motion across consecutive frames.
-To address it, a naive solution is to develop an agentic system that first leverages off-the-shelf subject-driven image generation models (Ruiz et al., 2023) to generate the first frame given the text prompt and reference images, and then applies video generative models (OpenAI, 2023; Yang et al., 2024b; Wan, 2025) to produce the subsequent frames.
-However, this two-stage pipeline lacks explicit correspondence between the video and the reference subjects, and heavily depends on the quality of the initial frame, resulting in suboptimal identity preservation, poor temporal consistency, and vulnerability to error accumulation.
-To tackle these challenges, recent works (Deng et al., 2025b; Liu et al., 2025b; Xue et al., 2025a) propose end-to-end solutions by injecting information of multiple subjects into the Multimodal Diffusion Transformers (MMDiT) (Esser et al., 2024) of pretrained video diffusions (Wan, 2025).
-Although effective, these methods simply concatenate visual tokens of all subjects and textual tokens to noise tokens.
-This results in higher perplexity during semantic integration and conflict resolution across different subjects, as well as suboptimal performance due to lacking explicit modeling of interactions among subjects and modalities.
-To this end, we propose ID-Composer, a novel framework for multi-subject video generation that achieves state-of-the-art performance in identity preservation, temporal consistency, and overall video quality.
-The key designs of ID-Composer are twofold: (1) A hierarchical identity-preserving attention mechanism.
-The task we are addressing involves multiple subjects and modalities, making it crucial to effectively aggregate features both within and across different conditions.
-Inspired by recent work using local-global attention (Wang et al., 2025a; Lin et al., 2025), we design a hierarchical attention mechanism that performs attention on three progressive levels: intra-subject, inter-subject, and cross-modal attention.
-This design enables effective identity consistency and textual faithfulness by capturing fine-grained details within each subject, modeling interactions among multiple subjects, and integrating information across modalities.
-(2) Semantic understanding via pretrained vision-language model (VLM).
-Our model leverages VLM as a text encoder to provide fine-grained guidance and capture complex interactions between multiple subjects.
-Compared to traditional text encoders (Radford et al., 2021), VLMs (Yang et al., 2024a) demonstrate superior semantic understanding and cross-modal alignment capabilities, which are crucial for the multi-subject video generation.
-ID-Composer is built upon a strong video diffusion model (Wan, 2025) and incorporates the above two designs into the MMDiT (Esser et al., 2024) architecture, enabling effective generation of high-quality videos with multiple specific subjects.
-To incentivize the model to generate videos with high quality and strong identity preservation, we further introduce an online reinforcement learning (Liu et al., 2025a; Guo et al., 2025) phase subsequent to the initial diffusion training.
-A new dataset is constructed to facilitate the training and evaluation of multi-subject video generation models.
-We introduce a data-curation pipeline leveraging vision-language models and advanced image editing models.
-Our process involves automated filtering, captioning, and subject-decomposing stages to process video clips, resulting in a mixed-modality dataset.
-As shown in Fig.
-1, ID-Composer is capable of generating high-quality videos that accurately preserve the identities of multiple subjects, align with the prompt, and maintain temporal consistency.
-Experiments demonstrate that our model surpasses existing methods in identity preservation, temporal consistency, and video quality.
-Ablation studies validate the effectiveness of our proposed designs.
-In summary, our contributions are as follows:
-  • We propose ID-Composer, a novel framework for multi-subject video generation that effectively preserves identities, integrates semantics, and maintains temporal consistency.
-  • Our proposed hierarchical identity-preserving attention mechanism, the use of pretrained VLM for semantic understanding, and the incorporation of online reinforcement learning significantly enhance the model’s performance in handling multiple subjects and modalities.
-  • We construct a new dataset for multi-subject video generation, and conduct extensive experiments demonstrating the superiority of ID-Composer over existing methods.
+Recent advances in video generative models [OpenAI2023Sora, yang2024cogvideox, Vidu, Keling, wan2025] have achieved remarkable fidelity video generation, yet their primary reliance on sparse inputs like a single text prompt or initial frame significantly limits their controllability, especially for complex scenes.
+Drawing inspiration from the compositional flexibility of modern image generators [deng2025emergingpropertiesunifiedmultimodal, nanobanana], we focus on the challenging task of multi-subject video generation.
+This capability is crucial for downstream applications such as subject-driven video synthesis, dynamic scene composition, and controllable product placement, promising to advance personalized content creation, virtual storytelling, and advertising.
+However, extending existent single-subject video generation techniques [he2024id, yuan2024identity, xue2025standinlightweightplugandplayidentity, sang2025lynx] to multiple subjects introduces a significant challenge: maintaining the distinct identity of each subject while ensuring temporal coherent throughout the video.
+Recent studies [CINEMA, liu2025phantom] attempt to address this by injecting multi-subject information into pretrained video diffusion models [wan2025].
+But these methods struggle to resolve semantic conflicts between subjects and leads to a degradation of identity.
+The fundamental challenge lies in the inherent tension between preserving individual subject identities and generating a coherent, dynamic scene—a conflict that current methods fail to resolve effectively.
+To address this dilemma, we propose ID-Crafter, which harmonizes identity consistency with motion quality by integrating three strategic components.
+ID-Crafter is specifically engineered to tackle the inherent challenges of multi-subject video synthesis: (1) First, to disentangle complex inter-subject relationships while preserving individual attributes, we introduce a hierarchical identity-preserving attention mechanism.
+This module operates in a cascaded manner: it first attends to intra-subject details to capture unique features, subsequently models inter-subject interactions to mitigate identity leakage, and finally performs cross-modal attention to ensure semantic alignment with the textual prompt.
+This hierarchical structure establishes a robust foundation for identity preservation.
+(2) Second, to enhance generation fidelity for complex descriptions, we utilize a foundational VLM [Qwen2.5-VL] for semantic feature extraction.
+unlike conventional text encoders [radford2021learning], VLMs offer a granular understanding of scene composition and cross-modal dynamics, which is essential for accurately interpreting and rendering intricate multi-subject scenarios.
+(3) Third, to directly optimize the trade-off between identity preservation, video quality, and motion fluidity, we introduce an online Reinforcement Learning (RL) strategy.
+Optimizing a multi-subject diffusion policy presents significant stability challenges.
+Inspired by Flow-GRPO [flowgrpo], we address this by employing groupwise comparison for advantage estimation [shao2024deepseekmath, Deepseek-R1], thereby eliminating the need for a fragile value network.
+Since applying Online RL (e.g., ReFL [ReFL] or DRaFT [DRaFT]) to a multi-subject generative architecture presents non-trivial challenges, as the optimization policy encompasses the video diffusion model augmented with our hierarchical ID attention.
+We address this by utilizing groupwise comparison for advantage estimation [shao2024deepseekmath, Deepseek-R1], thereby eliminating the reliance on a fragile value network.
+Furthermore, we introduce a contrastive learning mechanism that provides a stable training signal to fine-tune the hierarchical attention, directly maximizing identity fidelity while mitigating reward hacking.
+As illustrated in Fig.
+1, these synergistic components empower ID-Crafter to generate high-fidelity videos distinguished by precise identity preservation, robust prompt alignment, and exceptional temporal consistency.
+Consequently, our method markedly surpasses previous methods (e.g., Phantom [liu2025phantom]), in producing subject-consistent videos.
+Our contributions are summarized as follows:
+  • We propose ID-Crafter, a unified framework that mitigates the tension between identity preservation and motion fluidity in multi-subject video generation through the integration of hierarchical attention, VLM-based semantic guidance, and RL-based fine-tuning.
+  • We pioneer the application of online RL to video generation, demonstrating its effectiveness and stability in optimizing complex perceptual rewards and balancing competing generation objectives.
+  • Extensive experiments show the superiority of proposed method over existing methods in terms of identity preservation, temporal consistency, and overall video quality.
 2 Related Works
@@
 A prominent line of work involves tuning-based methods, which achieve high subject fidelity by fine-tuning a generative model on a small set of reference images.
-These approaches range from optimizing the entire model’s weights, as in DreamBooth (Ruiz et al., 2023), to learning subject-specific textual embeddings (Gal et al., 2022) or adopting parameter-efficient fine-tuning strategies (Hu et al., 2021; Han et al., 2023; Yuan et al., 2023).
+These approaches range from optimizing the entire model’s weights, as in DreamBooth [ruiz2023dreambooth], to learning subject-specific textual embeddings [gal2022image] or adopting parameter-efficient fine-tuning strategies [hu2021lora, han2023svdiff, yuan2023inserting].
 However, a significant drawback of these methods is the substantial computational overhead required for per-subject optimization.
-To circumvent this limitation, tuning-free methods (Ye et al., 2023; Wang et al., 2024; Sun et al., 2024; Wu et al., 2024; Liu et al., 2025c) have been proposed, which inject identity information during inference via various conditioning mechanisms, thereby offering a more efficient alternative.
-More recently, a new paradigm has emerged with versatile, unified models (Chen et al., 2024; Xiao et al., 2024; Lu, 2024; Wu et al., 2025a; Gao et al., 2025; Fortin et al., 2025) that incorporate subject-driven generation as one of many capabilities.
+To circumvent this limitation, tuning-free methods [ye2023ip, wang2024instantid, sun2024rectifid, wu2024infinite, liu2025mosaic] have been proposed, which inject identity information during inference via various conditioning mechanisms, thereby offering a more efficient alternative.
+More recently, a new paradigm has emerged with versatile, unified models [chen2024unireal, xiao2024omnigen, erwold-2024-qwen2vl-flux, wu2025qwen, seedream, nanobanana] that incorporate subject-driven generation as one of many capabilities.
 Despite this progress, extending subject-driven generation from the image to the video domain remains a formidable challenge, primarily due to the difficulty of maintaining both identity and temporal consistency across frames.
 Subject-Consistent Video Generation
-Subject-consistent video generation (keling, 2024; Vidu, 2024; Pika, 2024; Jiang et al., 2025) is commonly achieved by enhancing the model’s attention mechanisms to incorporate appearance cues.
-More flexible paradigms have also emerged, including adapter-based methods like ID-Animator (He et al., 2024), ConsisID (Yuan et al., 2024), and Stand-In (Xue et al., 2025a).
-For the specific challenge of handling multiple subjects, a router mechanism has been introduced (Mou et al., 2025).
-Other works have explored richer customization dimensions: Wang et al.
-(2025b) design FantasyPortrait to enhance multi-character portrait animation through expression-augmented diffusion transformers.
-Hu et al.
-(2025) propose PolyVivid to improve vividness and consistency in multi-subject scenarios via cross-modal interaction.
-Huang et al.
-(2025) present Videomage, which enables both multi-subject and motion customization in text-to-video diffusion.
-Our work is most closely related to Phantom (Liu et al., 2025b), Concat-ID (Zhong et al., 2025), SkyReels-A2 (Fei et al., 2025), and CINEMA (Deng et al., 2025b), which focus on attention-based feature injection.
+Subject-consistent video generation [Keling, Vidu, Pika, VACE] is commonly achieved by enhancing the model’s attention mechanisms to incorporate appearance cues.
+More flexible paradigms have also emerged, including adapter-based methods like ID-Animator [he2024id], CustomVideo [wang2024customvideo], ConsisID [yuan2024identity], and Stand-In [xue2025standinlightweightplugandplayidentity].
+For the specific challenge of handling multiple subjects, a router mechanism has been introduced [mou2025dreamo].
+Other works have explored richer customization dimensions: wang2025fantasyportrait design FantasyPortrait to enhance multi-character portrait animation through expression-augmented diffusion transformers.
+hu2025polyvivid propose PolyVivid to improve vividness and consistency in multi-subject scenarios via cross-modal interaction.
+huang2025videomage present Videomage, which enables both multi-subject and motion customization in text-to-video diffusion.
+Our work is most closely related to Phantom [liu2025phantom], Concat-ID [Concat-ID], SkyReels-A2 [fei2025skyreels], and CINEMA [CINEMA], which focus on attention-based feature injection.
 However, these methods, as well as the above extensions, often struggle with conflicts between subject identity and textual prompts and with maintaining consistency.
-Moreover, ID-Composer proposes a novel method to mitigate these conflicts and enforce multi-level consistency, thereby improving the robustness and quality of multi-subject video generation.
+Moreover, ID-Crafter proposes a novel method to mitigate these conflicts and enforce multi-level consistency, thereby improving the robustness and quality of multi-subject video generation.
 Reinforcement Learning for Generation
-Leveraging its computational efficiency, Direct Preference Optimization (DPO) (Gao et al., 2025) has become a prevalent technique for post-training text-to-image models to align them with human preferences, such as aesthetics and prompt fidelity.
-Subsequent work, like DenseDPO (Wu et al., 2025b), extends this paradigm to the video domain by refining paired data construction and the granularity of preference labels, yielding videos with superior visual quality and motion dynamics.
+Leveraging its computational efficiency, Direct Preference Optimization (DPO) [seedream] has become a prevalent technique for post-training text-to-image models to align them with human preferences, such as aesthetics and prompt fidelity.
+Subsequent work, like DenseDPO [DenseDPO], extends this paradigm to the video domain by refining paired data construction and the granularity of preference labels, yielding videos with superior visual quality and motion dynamics.
 A primary limitation of these methods, however, is their offline nature, which precludes online parameter updates.
-This challenge has been recently addressed by verifiable reward-based models like OpenAI’s o1 (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025).
-Concurrently, the growing interest in online learning has spurred various online Reinforcement Learning (RL) methods, including ReFL (Xu et al., 2023), DRaFT (Clark et al., 2023), FlowGRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025b), and SRPO (Shen et al., 2025), which have demonstrated promising results in enhancing image aesthetics and text-rendering fidelity.
+This challenge has been recently addressed by verifiable reward-based models like OpenAI’s o1 [OpenAI-O1] and Deepseek-R1 [Deepseek-R1, shao2024deepseekmath].
+Concurrently, the growing interest in online learning has spurred various online Reinforcement Learning (RL) methods, including ReFL [ReFL], RAFT [raft], DRaFT [DRaFT], FlowGRPO [flowgrpo], DanceGRPO [dancegrpo], and SRPO [SRPO], which have demonstrated promising results in enhancing image aesthetics and text-rendering fidelity.
 Nevertheless, their application to subject-consistent video generation remains limited.
 To bridge this gap, we introduce task-specific rewards tailored for subject-consistent video generation, a method that mitigates reward hacking and delivers videos with high identity consistency.
-Figure 2: ID-Composer Overview.
+Figure 2: ID-Crafter Overview.
 Our model incorporates a hierarchical identity-preserving attention mechanism and a VLM that performs reasoning on the multimodal input into a video DiT to enable multi-subject video generation.
@@
 3.4).
-3.1 Preliminaries: Rectified Flow for Video Generation
-Our methodology is founded upon a latent video diffusion transformer (Peebles & Xie, 2023; Wan, 2025) trained with Rectified Flow (RF) (Lipman et al., 2022; Liu et al., 2022; Esser et al., 2024).
+3.1 Rectified Flow for Video Generation
+Our methodology is founded upon a latent video diffusion transformer [peebles2023scalable, wan2025] trained with standard Rectified Flow (RF) [lipman2022flow, liu2022flow, esser2024scaling].
 In latent video diffusion, a video clip $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$ is first projected into a latent space by the encoder $\mathcal{E}$ of a variational autoencoder (VAE), yielding the representation $\mathbf{z}_{0}:=\mathcal{E}(\mathbf{V})\in\mathbb{R}^{f\times c\times h\times w}$.
@@
 The trajectory is parameterized by a time variable $t\in[0,1]$:
-$\mathbf{z}_{t}:=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}.$ (1)
+$\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}.$ (1)
 This path induces a constant velocity field $\mathbf{v}=\bm{\epsilon}-\mathbf{z}_{0}$.
@@
 The training objective is formulated as a regression loss over this velocity field:
-$\mathcal{L}_{\text{RF}}:=\mathbb{E}_{t,\mathbf{z}_{0},\bm{\epsilon}}\left[w(t)\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{C}_{\text{ctx}})-(\bm{\epsilon}-\mathbf{z}_{0})\right\rVert_{2}^{2}\right],$ (2)
+$\mathcal{L}_{\text{RF}}=\mathbb{E}_{t,\mathbf{z}_{0},\bm{\epsilon}}\left[w(t)\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{C}_{\text{ctx}})-(\bm{\epsilon}-\mathbf{z}_{0})\right\rVert_{2}^{2}\right],$ (2)
 where $\mathbf{C}_{\text{ctx}}$ represents the conditioning information (e.g., a text or image prompt), and $w(t)$ is a weighting function that balances the loss across different noise levels.
-3.2 ID-Composer: Multi-Subject Video Generation
+3.2 ID-Crafter: Multi-Subject Video Generation
 As shown in Fig.
-2, ID-Composer extends the DiT-based latent video diffusion architecture by introducing two key architectural innovations to handle the complexities of multi-subject conditioning: (1) A hierarchical identity-preserving attention mechanism, which is designed to hierarchically aggregate features.
-(2) Semantic understanding via a pretrained Vision-Language Model (VLM), which we employ as a sophisticated encoder for both textual and visual inputs.
+2, ID-Crafter extends the DiT-based latent video diffusion architecture by introducing two key architectural innovations to handle the complexities of multi-subject conditioning: (1) A hierarchical identity-preserving attention mechanism, which is designed to hierarchically aggregate features.
+(2) Semantic understanding via a pretrained VLM, which we employ as a sophisticated encoder for both textual and visual inputs.
 To further incentivize the model’s semantic alignment and identity preservation, we incorporate an online reinforcement learning phase subsequent to the initial flow matching training.
-3.2.1 Model Architecture
-Hierarchical Identity-Preserving Attention.
-Our model is based on multi-modal DiT (MMDiT), where latent video tokens and conditioning tokens are processed together through a series of transformer blocks.
-Following this, we first encode the input reference images $\mathcal{I}$ using a pretrained image encoder (e.g., Wan-VAE) to obtain visual feature maps $\{\mathbf{F}_{k}\}_{k=1}^{N}$, where each $\mathbf{F}_{k}\in\mathbb{R}^{c\times h\times w}$ corresponds to the $k$-th subject.
+Hierarchical Attention and Semantic Grounding.
+Our model is based on a Wan Video [wan2025], where latent video tokens and conditioning tokens are processed together.
+We first encode the input reference images $\mathcal{I}$ using a pretrained image encoder to obtain visual feature maps $\{\mathbf{F}_{k}\}_{k=1}^{N}$, where each $\mathbf{F}_{k}\in\mathbb{R}^{c\times h\times w}$ corresponds to the $k$-th subject.
 These feature maps are then flattened into token sequences $\{\mathbf{f}_{k}\}_{k=1}^{N}$, with each $\mathbf{f}_{k}\in\mathbb{R}^{hw\times c}$.
-The textual prompt $\mathbf{C}_{\text{txt}}$ is encoded using a pretrained text encoder to produce a sequence of text tokens $\mathbf{f}_{\text{txt}}\in\mathbb{R}^{l\times c}$, where $l$ is the number of text tokens.
-The complete conditioning token set is thus $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]$.
-To effectively integrate information from multiple subjects and modalities, we introduce a hierarchical attention mechanism within each transformer block.
-This mechanism operates in three stages: (1) Intra-subject attention: Each subject’s tokens $\mathbf{f}_{k}$ undergo self-attention independently, allowing the model to capture fine-grained details and spatial relationships within each subject.
-(2) Inter-subject attention: The outputs from the intra-subject attention are concatenated and subjected to another self-attention layer, enabling the model to learn interactions and relationships between different subjects.
-(3) Multi-modal attention: Finally, the combined subject features are concatenated with the text tokens $\mathbf{f}_{\text{txt}}$ and the latent video tokens, followed by a self-attention operation that fuses information across modalities.
-This hierarchical approach ensures that the model can effectively preserve individual subject identities while also capturing complex inter-subject dynamics and aligning them with the textual prompt.
-Semantic Reasoning via Vision-Language Models.
-While the hierarchical attention mechanism is designed to preserve identity, composing multiple subjects coherently in accordance with ambiguous user instructions remains a formidable challenge.
-Furthermore, models trained on limited data often lack the reasoning capabilities required for robust generalization to diverse real-world scenarios.
-To address these limitations, we leverage a powerful pretrained Vision-Language Model (VLM) to achieve a more sophisticated semantic understanding of the prompt and guide the composition of multiple subjects.
-We employ the Qwen2.5-VL model (Bai et al., 2025) for its superior multimodal reasoning capabilities.
-The VLM processes both the textual prompt $\mathbf{C}_{\text{txt}}$ and the set of reference images $\mathcal{I}$ to generate semantically enriched text tokens, which we denote as $\mathbf{f}_{\text{txt}}$: $\mathbf{f}_{\text{txt}}=\text{VLM}_{\text{enc}}(\mathbf{C}_{\text{txt}},\mathcal{I})\in\mathbb{R}^{l^{\prime}\times c}.$ These tokens capture not only the textual semantics but also the visual concepts of the subjects, enabling a more holistic understanding of the user’s intent.
-These VLM-enhanced tokens then replace the embeddings from a standard text encoder.
-The complete set of conditioning tokens $\mathbf{C}_{\text{ctx}}$ is subsequently formed by concatenating these enhanced text tokens with the individual subject tokens: $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]\in\mathbb{R}^{(l^{\prime}+N\cdot hw)\times c}.$ By integrating the VLM’s advanced reasoning, ID-Composer achieves superior semantic alignment and identity preservation, resulting in higher-quality videos that faithfully adhere to complex, multi-subject prompts.
-3.3 Training with Online Reinforcement Learning
-GRPO
-We use Flow-GRPO (Liu et al., 2025a) to perform online reinforcement learning (RL) to further improve the video generation quality.
-For each condition $q$, GRPO (Guo et al., 2025) samples a group of outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model by maximizing the following objective:
-$\mathcal{J}_{GRPO}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (3)
-$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{t}^{i}\left(\theta\right)\hat{A}_{i,t},\ \text{clip}\left(r_{t}^{i}\left(\theta\right),1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\,\|\,\pi_{\text{ref}}\big]\right\},$
-$\text{where}\quad r_{t}^{i}\left(\theta\right)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$
-the relative quality $\hat{A}_{i,t}$ of the i-th response is computed as
-$\hat{A}_{i,t}=\frac{r_{i}-\text{Mean}(\{r_{1},r_{2},\ldots,r_{n}\})}{\text{Std}(\{r_{1},r_{2},\ldots,r_{n}\})},$
-Flow-GRPO extends GRPO to the rectified flow setting, while convert flow matching’s determinsitic generation process to a stochastic one by injecting noise at each step of the flow integration.
-The update rule is defined as:
-$\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}+\left[{\bm{v}}_{\bm{\theta}}(\mathbf{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\left(\mathbf{x}_{t}+(1-t){\bm{v}}_{\theta}(\mathbf{x}_{t},t)\right)\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\bm{\epsilon},$ (4)
+To effectively integrate information from multiple subjects and modalities, we introduce a hierarchical identity-preserving attention mechanism.
+This operates in three stages: (1) Intra-subject attention captures fine-grained details within each subject; (2) Inter-subject attention models the interactions between different subjects; and (3) Multi-modal attention fuses subject features with text and video tokens.
+This hierarchical design is crucial for preserving individual identities while capturing complex inter-subject dynamics.
+Recently, VLMs have seen increasing adoption in modern video generation models [gao2025seedance, zhang2025waver], yet their potential for disentangling complex multi-subject dynamics remains significantly underutilized.
+Notably, no prior open-source work has integrated a VLM as the core reasoning engine within the powerful open-source Wan-Video architecture.
+Our approach fundamentally differs by harnessing the VLM’s granular cross-modal reasoning to actively guide our Hierarchical Identity-Preserving Attention.
+This transforms the VLM from a static encoder into a dynamic semantic guide.
+We employ the Qwen2.5-VL model [Qwen2.5-VL] to process the textual prompt $\mathbf{C}_{\text{txt}}$ and reference images $\mathcal{I}$, yielding semantically enriched tokens: $\mathbf{f}_{\text{txt}}=\text{VLM}_{\text{enc}}(\mathbf{C}_{\text{txt}},\mathcal{I})\in\mathbb{R}^{l^{\prime}\times c}$.
+This structured representation provides an explicit signal to the generative process.
+The complete set of conditioning tokens $\mathbf{C}_{\text{ctx}}$ is subsequently formed by concatenating these enhanced text tokens with the individual subject tokens: $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]\in\mathbb{R}^{(l^{\prime}+N\cdot hw)\times c}.$ Our central innovation is thus bridging the VLM’s high-level semantic reasoning with the generative model’s low-level spatial layout control.
+3.3 Post-training with Online RL
+Post-training the base model with a perceptual reward for identity preservation presents a formidable challenge, primarily due to the credit assignment problem.
+How to adjust the intricate computations within the hierarchical ID attention layers to enhance a holistic, video-level reward?
+While DPO-style methods [DenseDPO, DPO] offer computational efficiency, their offline nature precludes the online parameter updates essential for this fine-tuning task.
+Alternatively, standard policy-gradient algorithms [ReFL, DRaFT] are often impractical in this context.
+They necessitate training a separate value function to estimate rewards, which is unstable and sensitive to hyperparameters.
+Following recent advances [flowgrpo, dancegrpo], for each condition $q$, ID-Crafter samples a group of outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model by maximizing the following objective:
+$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{\begin{subarray}{c}q\sim P(Q),\\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)\end{subarray}}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}$ (3)
+$\quad\left\{\min\left[r_{t}^{i}(\theta)\hat{A}_{i,t},\text{clip}(r_{t}^{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]\right$.
+$\quad\left.-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\,\|\,\pi_{\text{ref}}\big]\right\},$
+$\text{where}\quad r_{t}^{i}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$
+the relative quality $\hat{A}_{i,t}$ of the i-th response is computed as $\hat{A}_{i,t}=\frac{r_{i}-\text{Mean}(\{r_{1},r_{2},\ldots,r_{n}\})}{\text{Std}(\{r_{1},r_{2},\ldots,r_{n}\})},$
+While convert flow matching’s determinsitic generation process to a stochastic one by injecting noise at each step of the flow integration. The update rule is defined as:
+$\mathbf{x}_{t+\Delta t}=\mathbf{x}_{t}$ $+\left[{\bm{v}}_{\bm{\theta}}(\mathbf{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\left(\mathbf{x}_{t}+(1-t){\bm{v}}_{\theta}(\mathbf{x}_{t},t)\right)\right]\Delta t$ (4)
+$+\sigma_{t}\sqrt{\Delta t}\,\bm{\epsilon}$
 where $\bm{\epsilon}\sim\mathcal{N}(0,I)$ injects stochasticity into the generation process and $\sigma_{t}=a\sqrt{\frac{t}{1-t}}$.
-Reward
-To further refine the model’s output, we propose an online reinforcement learning (RL) phase subsequent to the initial RF training.
-This stage employs a bespoke policy-gradient algorithm to fine-tune the generative model $\mathbf{v}_{\bm{\theta}}$.
-The objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of video quality and identity consistency metrics:
-$\bm{\theta}^{*}=\arg\max_{\bm{\theta}}\mathbb{E}_{\mathbf{V}\sim p_{\bm{\theta}}}\left[\mathcal{R}_{\text{total}}(\mathbf{V})\right],$ (5)
-where $\mathcal{R}_{\text{total}}(\mathbf{V}):=\lambda_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V})+\lambda_{\text{id}}\mathcal{R}_{\text{id}}(\mathbf{V},\mathcal{I})$.
-Here, $\mathcal{R}_{\text{qual}}$ assesses the perceptual quality of the video, while $\mathcal{R}_{\text{id}}$ quantifies the identity preservation with respect to the reference images $\mathcal{I}$.
-The weights $\lambda_{\text{qual}}$ and $\lambda_{\text{id}}$ balance these two objectives.
+Reward Design.
+The objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of video quality and identity consistency:
+$\mathcal{R}_{\text{total}}(\mathbf{V}):=w_{\text{fid}}\mathcal{R}_{\text{fid}}(\mathbf{V},\mathcal{I})+w_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V}),$ (5)
+where we empirically set the weights to $w_{\text{fid}}=0.6$, $w_{\text{qual}}=0.4$ to prioritize identity preservation while ensuring high quality and text alignment.
+Fidelity Reward ($\mathcal{R}_{\text{fid}}$).
+The fidelity reward ensures that the generated subjects are consistent with the reference images $\mathcal{I}$. It combines a face-specific metric with a holistic subject consistency score:
+$\mathcal{R}_{\text{fid}}:=(1-\alpha)\mathcal{R}_{\text{face}}+\alpha\mathcal{R}_{\text{subject}}.$ (6)
+For facial identity $\mathcal{R}_{\text{face}}$, We use a robust combination of the mean and minimum FaceSim scores [consisid] across all $N$ subjects:
+$\mathcal{R}_{\text{face}}:=(1-\gamma)\left(\frac{1}{N}\sum_{k=1}^{N}\mathcal{R}_{\text{id}}^{k}\right)+\gamma\left(\min_{k=1,\dots,N}\mathcal{R}_{\text{id}}^{k}\right),$ (7)
+where $\mathcal{R}_{\text{id}}^{k}$ is the ArcFace-based score for subject $k$, and we set $\gamma=0.5$.
+For overall subject consistency, $\mathcal{R}_{\text{subject}}$, we employ the NexusScore, a metric from our proposed OpenS2V-Eval benchmark that assesses non-facial and holistic subject attributes.
+We balance face and subject with $\alpha=0.5$.
+Quality Reward ($\mathcal{R}_{\text{qual}}$).
+The quality reward assesses both the aesthetic appeal and the physical plausibility of the generated video $\mathbf{V}$:
+$\mathcal{R}_{\text{qual}}:=(1-\beta)\mathcal{R}_{\text{aes}}+\beta\mathcal{R}_{\text{nat}}.$ (8)
+$\mathcal{R}_{\text{aes}}$ is the standard Aesthetic Score [improved-aesthetic-predictor].
+To combat reward hacking that may produce unnatural but aesthetically pleasing artifacts, we introduce $\mathcal{R}_{\text{nat}}$.
+This is the NaturalScore from our benchmark, which leverages a powerful VLM to evaluate whether the video adheres to physical laws and common sense.
+We set $\beta=0.4$ to strongly penalize unnatural generations.
+Figure 3: Data curation pipeline of ID-Crafter.
 3.4 Training Data curation
-Figure 3: Data curation pipeline of ID-Composer.
 Current subject-to-video generation methods are hampered by notable deficiencies in output quality and diversity.
 These shortcomings are largely attributable to the restrictive nature of paired training data, which proves inadequate for modeling complex, real-world variations in subject motion, camera viewpoint, and scene layout.
-To surmount these limitations, we propose a sophisticated data-curation pipeline that harnesses the capabilities of modern Vision-Language Models (VLMs) (Bai et al., 2025) and powerful image editing models (Fortin et al., 2025).
+To surmount these limitations, we propose a sophisticated data-curation pipeline that harnesses the capabilities of modern VLMs (e.g., QwenVL-72B) [Qwen2.5-VL] and a powerful image editing model (a.k.a., Nano Banana) [nanobanana].
+The first component consists of subject-video pairs extracted from a diverse collection of OpenS2V-Nexus [yuan2025opens2v], offering a wide array of authentic scenes and actions.
+The second component comprises synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, including explicitly designed cross-subject compositions and fusion examples, to mitigate “copy-plate” artifacts and enhance multi-entity interaction coherence.
+The third component is a high-fidelity collection of professionally shot videos with detailed annotations.
 As shown in Fig.
@@
 This dataset is composed of three heterogeneous data sources.
-The first component consists of subject-video pairs extracted from a diverse collection of OpenS2V-Nexus (Yuan et al., 2025), offering a wide array of authentic scenes and actions.
-The second component comprises synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, systematically increasing the diversity of subject-background compositions.
-The third component is a high-fidelity collection of professionally shot videos with detailed annotations.
-More technical details are provided in Appendix C.
+More technical details are provided in Appendix.
 4 Experiments
 4.1 Experimental Settings
+[…]
+Proprietary Models
+Vidu 2.0 [Vidu] 47.59% 41.47% 13.52% 35.11% 67.57% 43.55% 71.44%
+Pika 2.1 [Pika] 48.88% 46.87% 24.70% 30.80% 69.21% 45.41% 69.79%
+Kling 1.6 [Keling] 54.46% 44.60% 41.60% 40.10% 66.20% 45.92% 79.06%
+Open-source Models
+VACE-1.3B [VACE] 45.53% 48.24% 18.83% 20.58% 71.26% 37.95% 71.78%
+VACE-14B [VACE] 52.87% 47.21% 15.02% 55.09% 73.27% 44.20% 72.78%
+Phantom-1.3B [liu2025phantom] 50.71% 46.67% 14.29% 48.55% 69.43% 42.44% 70.26%
+Phantom-14B [liu2025phantom] 52.32% 46.39% 33.42% 51.48% 70.65% 37.43% 68.66%
+SkyReels-A2-P14B [fei2025skyreels] 49.61% 39.40% 25.60% 45.95% 64.54% 43.77% 67.22%
+Ours-1.3B (Base Model)
+$\star$
+54.33% 42.50% 38.00% 58.12% 64.06 % 43.22% 71.09%
+Ours-1.3B (Base Model + online RL)
+$\star$
+55.16% 48.85% 36.50% 66.10% 62.20% 43.45% 69.15%
+Ours-14B (Large-scale Model) 57.05% 45.28% 40.34% 60.71% 67.49% 45.11% 73.23%
+Table 1: Quantitative Comparison against existing methods for the open-domain subject-to-video benchmark.
+Total score is the normalized weighted sum of other scores.
+$\uparrow$” indicates that higher is better.
 Implementation Details
-Our model, ID-Composer, is initialized from the weights of the Wan-Video-1.3 B model (Wan, 2025).
-To leverage superior semantic understanding, we employ a dual text-encoder architecture, combining T5 (Raffel et al., 2020) with Qwen2.5-VL-7B-Instruct (Bai et al., 2025).
+Our model, ID-Crafter, is initialized from the weights of the Wan-Video-1.3 B model [wan2025].
+To leverage superior semantic understanding, we employ a dual text-encoder architecture, combining T5 [T5] with Qwen2.5-VL-7B-Instruct [Qwen2.5-VL].
 The model was trained for 30,000 iterations on our curated dataset at a resolution of 480p, utilizing 16 H20 GPUs.
-For inference, we employ an Euler sampler with 50 steps and classifier-free guidance (Ho & Salimans, 2022) to modulate the influence of image and text conditions, setting the classifier-free guidance scale to 2.5.
+For inference, we employ an Euler sampler with 50 steps and classifier-free guidance [ho2022classifier] to modulate the influence of image and text conditions, setting the classifier-free guidance scale to 2.5.
+The inference time for our 1.3B model is approximately one minute to generate a video at 480p resolution.
 Baselines
 We benchmark our method against state-of-the-art open-source models that support native subject-to-video generation.
-This includes Phantom (Liu et al., 2025b) and SkyReels-A2 (Fei et al., 2025), both available in 1.3B and 14B parameter variants, and VACE (Jiang et al., 2025).
-For a comprehensive assessment, we also present a qualitative comparison with leading proprietary systems, such as VIDU (Vidu, 2024), Pika (Pika, 2024) and Kling (keling, 2024).
+This includes Phantom [liu2025phantom] and SkyReels-A2 [fei2025skyreels], both available in 1.3B and 14B parameter variants, and VACE [VACE].
+For a comprehensive assessment, we also present a qualitative comparison with leading proprietary systems, such as VIDU [Vidu], Pika [Pika] and Kling [Keling].
 Evaluation Metrics
-Following the benchmark protocol of OpenS2V-Nexus (Yuan et al., 2025), we perform evaluation on a diverse dataset of 180 unique subject-text pairs.
+Following the benchmark protocol of OpenS2V-Nexus [yuan2025opens2v], we perform evaluation on a diverse dataset of 180 unique subject-text pairs.
+To ensure a rigorous and unbiased evaluation, this dataset is maintained as a held-out set, completely disjoint from our training data, thereby preventing any potential data leakage.
 We report on a suite of automated metrics, including Aesthetics, Motion quality, Face Similarity (FaceSim), and alignment scores against three benchmarks (GmeScore, NexusScore, NaturalScore).
-The Total Score is a normalized weighted sum of these metrics.
-Further details are provided in Appendix D.
+In our ablation studies, we also report the VLM-based visual scorer, Q-Align [wu2023qalign], for perceptual quality.
 4.2 Main Results
 Quantitative Comparison
-As demonstrated in Table 1, ID-Composer achieves state-of-the-art performance across all evaluation metrics, outperforming existing open-source models by a significant margin.
+As demonstrated in Table 1, ID-Crafter achieves state-of-the-art performance across all evaluation metrics, outperforming existing open-source models by a significant margin.
 Notably, our method shows substantial improvements in FaceSim and NexusScore, which we attribute to our hierarchical identity-preserving attention mechanism and the rich semantic guidance from the VLM encoder.
 These components work in concert to ensure high-fidelity identity preservation and robust alignment with complex textual prompts.
-Table 1: Quantitative Comparison against existing methods for the open-domain subject-to-video benchmark. Total score is the normalized weighted sum of other scores. $\uparrow$” higher is better.
-[…]
-Proprietary Models
-Vidu 2.0 (Vidu, 2024) 47.59% 41.47% 13.52% 35.11% 67.57% 43.55% 71.44%
-Pika 2.1 (Pika, 2024) 48.88% 46.87% 24.70% 30.80% 69.21% 45.41% 69.79%
-Kling 1.6 (keling, 2024) 54.46% 44.60% 41.60% 40.10% 66.20% 45.92% 79.06%
-Open-source Models
-VACE-1.3B (Jiang et al., 2025) 45.53% 48.24% 18.83% 20.58% 71.26% 37.95% 71.78%
-VACE-14B (Jiang et al., 2025) 52.87% 47.21% 15.02% 55.09% 67.27% 44.20% 72.78%
-Phantom-1.3B (Liu et al., 2025b) 50.71% 46.67% 14.29% 48.55% 69.43% 42.44% 70.26%
-Phantom-14B (Liu et al., 2025b) 52.32% 46.39% 33.42% 51.48% 70.65% 37.43% 68.66%
-SkyReels-A2-P14B (Fei et al., 2025) 49.61% 39.40% 25.60% 45.95% 64.54% 43.77% 67.22%
-Ours-1.3B (Base Model) 54.33% 42.50% 38.00% 58.12% 64.00% 43.00% 71.00%
-Ours-14B 57.05% 45.00% 40.00% 60.50% 67.00% 45.00% 73.00%
 Qualitative Comparison
-Figure 4 presents a qualitative comparison against leading methods.
-ID-Composer consistently generates videos with superior visual quality, temporal coherence, and identity preservation.
+Fig.
+4 presents a qualitative comparison against leading methods.
+ID-Crafter consistently generates videos with superior visual quality, temporal coherence, and identity preservation.
 While baseline methods often exhibit artifacts or identity drift, especially in longer videos or complex scenes, our model maintains a stable representation of all subjects , accurately reflecting their appearance from the reference images.
 Figure 4: Qualitative Comparison with State-of-the-Art Methods.
-Our method, ID-Composer, demonstrates superior performance in identity preservation, temporal consistency, and alignment with the textual prompt compared to existing open-source and proprietary models.
+Our method, ID-Crafter, demonstrates superior performance in identity preservation, temporal consistency, and alignment with the textual prompt compared to existing open-source and proprietary models.
 4.3 Ablation and Analysis
-We conduct a series of ablation studies to validate the effectiveness of our key design choices. The results, summarized in Table 2, underscore the importance of each proposed component.
-Table 2: Ablation studies of ID-Composer.
-We evaluate the impact of our key contributions: the hierarchical identity-preserving attention, the VLM for semantic understanding, and our curated dataset.
-All metrics are reported on our evaluation set.
-$\uparrow$” indicates that higher is better.
-[…]
-ID-Composer (Full Model) 58.12% 49.55% 48.91% 54.33%
-w/o Hierarchical Attention 51.34% 48.98% 47.52% 50.11%
-w/o VLM Encoder 56.98% 42.17% 46.88% 49.89%
-w/o Curated Data 54.55% 45.32% 45.13% 48.78%
 Effect of Hierarchical ID-Preserving Attention
@@
 Effect of VLM-based Semantic Understanding
-We evaluated the impact of the VLM by replacing the dual-encoder setup with a single T5 encoder (Raffel et al., 2020).
+We evaluated the impact of the VLM by replacing the dual-encoder setup with a single T5 encoder [T5].
 The results (“w/o VLM Encoder”) show a marked decrease in Text-Video Alignment.
@@
 Finally, we trained the full model on a baseline dataset without our curated data.
-This version (”w/o Curated Data”) yielded lower scores across all metrics, particularly in Video Quality.
+This version (“w/o Curated Data”) yielded lower scores across all metrics, particularly in Video Quality.
+Critically, the base model exhibited more severe “copy-plate” artifacts, due to limited exposure to cross-subject interactions.
+Our data curation pipeline explicitly mitigates this by incorporating synthetic cross-subject compositions and fusion examples, which enhances the model’s ability to synthesize coherent multi-entity interactions.
 This demonstrates that our data curation pipeline, which synthesizes diverse and high-quality training examples, is vital for achieving robust and generalizable performance.
-Qualitative results in Figure 5 further illustrate these findings.
+Qualitative results in Fig.
+5 further illustrate these findings.
+Table 2: Ablation studies on the key components of ID-Crafter. We compare the full model against variants removing hierarchical attention, the VLM encoder, and the curated dataset.
+Method FaceSim $\uparrow$ Q-Align $\uparrow$ Video Quality $\uparrow$ Total Score $\uparrow$
+[…]
+$\star$
+58.12% 0.351 48.91% 54.33%
+   w/o Hierarchical Attention 51.34% ($\downarrow$11.7%) 0.348 ($\downarrow$0.9%) 47.52% ($\downarrow$2.8%) 50.11% ($\downarrow$7.8%)
+   w/o VLM Encoder 56.98% ($\downarrow$2.0%) 0.287 ($\downarrow$18.2%) 46.88% ($\downarrow$4.2%) 49.89% ($\downarrow$8.2%)
+   w/o Curated Data 54.55% ($\downarrow$6.1%) 0.321 ($\downarrow$8.5%) 45.13% ($\downarrow$7.7%) 48.78% ($\downarrow$10.2%)
 Figure 5: Qualitative Ablation Study.
-The left panel highlights the importance of our curated dataset, showing improved coherence and realism in subject integration compared to a model trained without it.
-The right panel illustrates the effectiveness of our online reinforcement learning stage, which significantly enhances visual quality and subject consistency.
-4.4 Application: Controllable Video Editing
-The architecture of ID-Composer naturally lends itself to a range of editing applications.
+The left panel highlights the importance of our curated dataset, demonstrating improved coherence and realism in subject integration (e.g., mitigating ‘copy-paste’ artifacts) compared to a model trained without it.
+Meanwhile, the right panel illustrates the effectiveness of our online reinforcement learning stage, which significantly enhances visual quality and subject consistency.
+Figure 6: Applications in Controllable Video Editing.
+ID-Crafter enables zero-shot editing of existing videos, including subject replacement and background modification, while preserving identity and temporal consistency.
+Effect of Online RL
+As shown in Tab.
+3, the online RL stage brings substantial improvements across all metrics compared to the SFT baseline.
+Specifically, our model achieves a 13.7% relative increase in FaceSim and a 14.9% relative increase in Aesthetics.
+The alignment score, Q-Align, also sees a significant 16.8% relative improvement.
+This demonstrates that directly optimizing the model with our composite reward function is crucial for enhancing identity preservation and perceptual quality.
+The qualitative results in Fig.
+5 (right panel) further corroborate this finding, where the model trained with RL generates videos with significantly better subject consistency compared to the SFT baseline.
+Table 3: Analysis of the online RL optimization.
+We compare our online GRPO approach with offline DPO and a standard SFT baseline.
+We also ablate the components of our composite reward function $\mathcal{R}_{\text{total}}$.
+Method FaceSim $\uparrow$ Aesthetics $\uparrow$ Q-Align. $\uparrow$ Total Score $\uparrow$
+SFT Baseline 58.12% 42.50% 0.351 54.33%
+Comparison of RL Algorithms
+DPO (Offline) 62.35% 45.15% 0.382 54.80%
+Ours (Online) 66.10% 48.85% 0.410 55.16%
+Ablation on Reward Components
+Ours w/o Fidelity ($\mathcal{R}_{\text{fid}}$) 45.32% 47.10% 0.391 53.50%
+Ours w/o Quality ($\mathcal{R}_{\text{qual}}$) 63.50% 43.81% 0.379 54.82%
+Ours w/o Natural ($\mathcal{R}_{\text{nat}}$) 69.30% 50.83% 0.361 53.01%
+Ours (Online) 66.10% 48.85% 0.410 55.16%
+Analysis of Online RL Optimization
+Offline RL methods like DPO [DPO] offer an alternative but rely on a static dataset of preference pairs, which may not sufficiently cover the vast space of possible generations.
+In contrast, our online approach allows the policy to actively explore and receive feedback from our reward model, leading to more effective optimization.
+As shown in Tab.
+LABEL:↑:rl_ablation, our online GRPO method significantly outperforms both the SFT baseline and the offline approach across all metrics.
+This confirms that online sample generation is more effective for fine-tuning video generation models.
+The table also presents an ablation study on the components of our composite reward function, $\mathcal{R}_{\text{total}}$.
+Removing the fidelity term ($\mathcal{R}_{\text{fid}}$) causes a severe drop in FaceSim, while removing the quality ($\mathcal{R}_{\text{qual}}$) term leads to a degradation in Aesthetics.
+Interestingly, removing the Natural Reward $\mathcal{R}_{\text{nat}}$ leads to a degradation in Q-Align, but a slight increase in FaceSim and Aesthetics.
+This suggests a reward hacking scenario, where the model over-optimizes for certain metrics at the expense of overall alignment.
+These results validate our multi-faceted reward design and the necessity of each component for achieving a balanced improvement.
+Controllable Video Editing
+The architecture of ID-Crafter naturally lends itself to a range of editing applications.
 By providing an existing video as an initial state and conditioning the generation on new subjects or modified text prompts, our model can perform zero-shot, identity-consistent video editing.
-As illustrated in Figure 6, ID-Composer can seamlessly insert new subjects into a scene, replace existing ones, or alter the background, all while maintaining temporal and semantic coherence.
+As illustrated in Fig.
+6, ID-Crafter can seamlessly insert new subjects into a scene, replace existing ones, or alter the background, all while maintaining temporal and semantic coherence.
 This capability opens up new avenues for personalized content creation and virtual storytelling.
-Figure 6: Applications in Controllable Video Editing.
-ID-Composer enables zero-shot editing of existing videos, including subject replacement and background modification, while preserving identity and temporal consistency.
 5 Conclusion
-In this work, we introduced ID-Composer, an innovative framework for multi-subject video generation that integrates hierarchical identity-preserving attention and a pretrained Vision-Language Model into a latent diffusion backbone.
-Moving beyond simple feature concatenation, ID-Composer captures complex intra- and inter-subject interactions, preserves identities, incorporates semantics, and ensures temporal consistency.
-Extensive experiments demonstrate that ID-Composer achieves state-of-the-art performance, outperforming existing methods across a comprehensive suite of metrics.
+In this work, we introduced ID-Crafter, an innovative framework for multi-subject video generation that integrates hierarchical identity-preserving attention and a VLM into a latent diffusion backbone.
+Moving beyond simple feature concatenation, ID-Crafter captures complex intra- and inter-subject interactions, preserves identities, incorporates semantics, and ensures temporal consistency.
+Extensive experiments demonstrate that ID-Crafter achieves state-of-the-art performance, outperforming existing methods across comprehensive metrics.
 Our work marks an important step towards general and scalable multi-subject video generation in complex, real-world scenarios.
-Limitations and Future Work
-While ID-Composer excels at identity preservation and multi-subject video generation, it remains limited by high computational costs and difficulties with complex interactions and fine-grained dynamics.
-Future work includes developing more efficient architectures, incorporating physics-aware priors and 3D reasoning, mitigating biases in pretrained components, and advancing controllable generation for fine-grained attribute, action, and interaction specification.
-Appendix A Declaration of LLM Usage
-During the writing of the manuscript, we utilized a Large Language Model (ChatGPT) as a writing assistant.
-The scope of its usage was limited to improving grammar, polishing sentences, and enhancing the clarity and fluency of this manuscript.
-The method, claims, experimental results and conclusions are developed by the authors.
-Appendix B More Implementation Details
-Reproducibility
-To facilitate reproducibility, we present our detailed experimental settings and evaluation metrics in Section 4.1.
-This section provides a comprehensive description of our implementation details.
-Moreover, our source code and pre-trained models will be publicly available.
-B.1 Details of Network
-We use a pre-trained VAE model from Wan-Video (Wan, 2025) to convert images from pixel space to latent space and vice versa.
-The latent representation has a down-sample ration of 8 and a latent channel of 16, and is then processed by a 2 × 2 patch embedding layer to reduce the spatial size and match the hidden dimension of the LLM backbone.
-The VAE model is frozen during training.
-Network Architecture
-Our video generation model is built upon three core modules: a Vision-Language Model (VLM) (Bai et al., 2025) for multi-modal comprehension, a Variational Autoencoder (VAE) for spatial compression, and a Multi-Modal Diffusion Transformer (DiT) (Wan, 2025) serving as the latent video diffusion backbone.
-The VLM, with a total of 7B parameters, is responsible for encoding textual prompts and subject images into a shared semantic space.
-It consists of a 32-layer Vision Transformer (ViT) and a 28-layer Large Language Model (LLM).
-The VAE is designed to encode video frames into a compact latent representation and subsequently decode them back to the pixel space.
-Its encoder and decoder utilize a spatial scale factor of $8\times 8$, managing a parameter count of 54M and 73M, respectively.
-The core of our generative process is the DiT, a substantial 14B parameter transformer that operates on the latent codes produced by the VAE.
-With 60 layers and an intermediate size of 12,288, it is architected to model complex spatio-temporal dynamics for high-fidelity video synthesis.
-A detailed breakdown of the architectural configurations for each component is provided in Table 3.
-Table 3: Detailed configuration of our model’s primary architectural components.
-The VLM, VAE, and MMDiT modules are designed to handle multimodal understanding, spatial compression, and latent diffusion, respectively.
-Configuration VLM VAE MMDiT
+Limitations and Future Work Despite its strength in multi-subject identity preservation, ID-Crafter has limitations in modeling complex interactions and fine-grained dynamics.
+Future work will focus on incorporating physics-aware priors, mitigating biases in pretrained components, and advancing fine-grained controllable generation of attributes, actions, and interactions.
+[…]
+Supplementary Material
+Appendices Overview
+The supplementary material includes the following sections:
+  • Sec. A: More Implementation Details.
+    • Details of Network Architecture (Sec. A.1).
+    • Details of Hierarchical Identity-Preserving Attention (Sec. A.1).
+    • Details of Reward Calculation (Sec. A.2).
+  • Sec. B: Dataset Curation Details.
+    • Dataset Composition and Statistics (Sec. B.1).
+    • Reference Image Construction Strategies (Sec. B.2).
+  • Sec. C: Evaluation Metric Details.
+  • Sec. D: Additional Results and Analysis.
+    • Evaluation on Multi-Subject Scenarios (Sec. D.1).
+    • Human Preference Study (Sec. D.2).
+  • Sec. E: Additional Discussion.
+    • Computational Overhead (Sec. E.1).
+    • Ethics Statement (Sec. E.2).
+Appendix A More Implementation Details
+Table S.4: Detailed configuration of our model’s primary architectural components.
+The VLM, VAE, and DiT modules are designed to handle multimodal understanding, spatial compression, and latent diffusion, respectively.
+Configuration VLM VAE DiT
 ViT LLM Encoder Decoder
-# Layers 32 28 11 15 60
-# Num Heads (Q / KV) 16 / 16 28 / 4 - - 24 / 24
-Head Size 80 128 - - 128
-Intermediate Size 3,456 18,944 - - 12,288
-Patch / Scale Factor 14 - $\times8$ 8 $\times$ 8 2
+# Layers 32 28 11 15 30/40
+# Num Heads (Q / KV) 16 / 16 28 / 4 - - 12 / 40
+Head Size 80 128 - - 128/128
+Intermediate Size 3,456 18,944 - - 8,960/13,824
+Patch / Scale Factor 14 - $\times8$ $\times8$ -
 Channel Size - - 16 16 -
-# Parameters 7B 54M 73M 1.3B / 14B
-B.2 Details of Reward Calculation
-Appendix C Dataset Curation Details
-C.1 Dataset Statistics
-Figure 7 illustrates the composition and distribution of scenarios in our training dataset, which is meticulously curated from three primary sources to ensure comprehensive coverage and high fidelity.
-The foundational component is the large-scale OpenS2V dataset (Yuan et al., 2025), which contributes a substantial volume of 218,230 videos, accompanied by 535,259 masked subject images and 3,098 generated counterparts.
-To further elevate the dataset’s quality, we incorporate a high-quality collection, comprising 9,374 pristine videos and an additional 3,000 generated subject images.
+[…]
+Table S.5: Ablation of Hierarchical Attention.
+We analyze the contribution of each stage in our hierarchical attention mechanism.
+The full three-stage model provides the best performance, with Stage 3 (cross-modal attention to VLM) being the most critical component for semantic alignment.
+Method FaceSim $\uparrow$ NexusScore $\uparrow$ Total Score $\uparrow$
+[…]
+w/o Stage 1 (Intra-Subject) 63.2% 44.2% 54.1%
+w/o Stage 2 (Inter-Subject) 64.5% 44.5% 54.5%
+w/o Stage 3 (Cross-Modal) 58.9% 38.7% 50.8%
+Vanilla Attention (Stage 3 only) 55.1% 39.1% 49.7%
+A.1 Network Architecture
+Our model’s architecture is founded on three synergistic modules: a Vision-Language Model (VLM), specifically the Qwen2.5-VL 7B variant [Qwen2.5-VL], for advanced multi-modal comprehension; a Variational Autoencoder (VAE) [rombach2022high] for efficient spatial compression; and a Diffusion Transformer (DiT) [peebles2023scalable, wan2025], which serves as the latent video diffusion backbone.
+The architectural configurations for each component are detailed in Table S.4.
+The 7B-parameter VLM, comprising a 32-layer Vision Transformer (ViT) and a 28-layer Large Language Model (LLM), is responsible for encoding textual prompts and subject images into a unified, semantically rich space.
+The VAE, with a 54M-parameter encoder and a 73M-parameter decoder, utilizes an $8\times$ spatial downsampling factor, compressing video frames into a compact latent representation that is computationally tractable.
+Subject Image Encoder
+We employ the pre-trained VAE from Wan-Video [wan2025] to transform subject images from the pixel space into the latent space.
+This process yields a $16$-channel latent representation with an $8\times$ downsampling factor.
+This latent representation is subsequently processed by a $2\times 2$ patch embedding layer, which further reduces its spatial dimensions to align with the hidden dimension of the DiT backbone, ensuring seamless integration between the visual and generative components.
+The VAE weights are kept frozen during training to preserve its high-fidelity reconstruction capabilities.
+Hierarchical Identity-Preserving Attention
+As detailed in the main paper, our Hierarchical Identity-Preserving Attention module is central to preventing identity leakage in multi-subject generation.
+It comprises three sequential stages applied within each block of the DiT, designed to progressively refine and integrate identity features.
+Let $Z_{l}$ be the video latent tokens at layer $l$, and $F_{subj}=\{f_{1},f_{2},\dots,f_{N}\}$ be the token sets for the $N$ subject images.
+Stage 1: Intra-Subject Attention.
+To refine the feature representation for each subject independently, we first apply a standard self-attention layer [vaswani2017attention] to each subject’s token set $f_{i}$.
+This initial stage allows the model to capture fine-grained, identity-specific details and consolidate features for each subject before any cross-subject interaction occurs.
+$f^{\prime}_{i}=\text{SelfAttention}(f_{i})\qquad\forall i\in\{1,\dots,N\}$ (S.9)
+Stage 2: Gated Inter-Subject Attention.
+To model interactions between subjects while strictly mitigating identity bleed, we employ a novel gated cross-attention mechanism.
+For each subject $f^{\prime}_{i}$, it queries the features of all other subjects $\{f^{\prime}_{j}\}_{j\neq i}$, which then modulates the output of the cross-attention.
+$f^{\prime\prime}_{i}=f^{\prime}_{i}+\sigma(W_{g}f^{\prime}_{i})\odot\text{CrossAttention}(Q=f^{\prime}_{i},KV=\{f^{\prime}_{j}\}_{j\neq i})$ (S.10)
+Stage 3: Cross-Modal Attention.
+Finally, the refined subject tokens $F^{\prime\prime}_{subj}$ and the textual context tokens $f_{txt}$ attend to the main video latent tokens $Z_{l}$.
+This stage injects the rich semantic guidance from the VLM, now conditioned on the robust and disentangled subject identities.
+While the VLM features [Qwen2.5-VL] are computed only once, their influence is dynamic, as they modulate the video features at each layer of the DiT.
+$Z^{\prime}_{l}=Z_{l}+\text{CrossAttention}(Q=Z_{l},KV=[F^{\prime\prime}_{subj};f_{txt}])$ (S.11)
+Table S.5 provides a detailed ablation study on this design.
+These three stages form a single, cohesive block that is inserted within the first 20 layers of the DiT backbone.
+This design choice is based on the observation that identity features are most critical in the early stages of the generation process.
+A.2 Details of GRPO-based Post-training
+This section provides a more detailed breakdown of the reward calculation and hyperparameter settings for our online post-training phase, which leverages the Groupwise Policy Optimization (GRPO) algorithm [flowgrpo].
+Advantage Normalization
+To stabilize the training process and reduce variance in the policy gradient updates, we normalize the advantage term.
+The normalized advantage $\hat{A}_{i}$ for the $i$-th sample in a group of size $G$ is calculated from the total trajectory reward $r_{i}$:
+$\hat{A}_{i}=\frac{r_{i}-\text{Mean}(\{r_{j}\}_{j=1}^{G})}{\text{Std}(\{r_{j}\}_{j=1}^{G})+10^{-8}}$ (S.12)
+where the total reward $r_{i}$ is computed by averaging the frame-level rewards over the video’s duration to produce a single scalar value.
+The small epsilon ($10^{-8}$) is added for numerical stability.
+The constant $a$ for noise injection $\sigma_{t}$ was set to 1.0.
+Reward Design
+As defined in the main paper, the objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of fidelity and quality rewards:
+$\mathcal{R}_{\text{total}}(\mathbf{V}):=w_{\text{fid}}\mathcal{R}_{\text{fid}}(\mathbf{V},\mathcal{I})+w_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V})$ (S.13)
+We empirically set the weights to $w_{\text{fid}}=0.6$ and $w_{\text{qual}}=0.4$ to prioritize identity preservation while ensuring high quality.
+The fidelity reward $\mathcal{R}_{\text{fid}}$ is a combination of face-specific and holistic subject consistency scores, while the quality reward $\mathcal{R}_{\text{qual}}$ assesses aesthetic appeal and physical plausibility.
+Hyperparameters and Sensitivity
+The key hyperparameters used during the GRPO training are detailed in Table S.6.
+We also conducted a sensitivity analysis on critical parameters, with results presented in Table S.7, to demonstrate the robustness of our training setup.
+Table S.6: Reinforcement Learning Hyperparameters. Key hyperparameters used for the online optimization stage with our GRPO algorithm.
+Hyperparameter Value
+Algorithm Flow-GRPO [flowgrpo, Deepseek-R1, shao2024deepseekmath]
+Policy Learning Rate 1e-5
+Optimizer AdamW
+[…]
+AdamW $\epsilon$ 1e-8
+Batch Size 32
+Max Gradient Norm 1.0
+Training Dynamics and Reward Effectiveness
+The effectiveness of our multi-faceted reward model is visualized in Figure S.7, which illustrates the evolution of key metrics during the online GRPO phase.
+The plot clearly shows a positive trend for identity fidelity (e.g., FaceSim), perceptual quality (Aesthetics), and Q-Align [wu2023qalign] over the course of training.
+This concurrent improvement across diverse axes validates that our reward model successfully guides the generator towards producing videos that are not only more identity-consistent but also more visually appealing and semantically accurate, without suffering from catastrophic forgetting or reward over-optimization on a single metric.
+Ablation on Fidelity Reward
+A crucial component of our total reward is the fidelity term $\mathcal{R}_{\text{fid}}$, which contains a balance factor $\gamma$ for the face-specific reward $\mathcal{R}_{\text{face}}$.
+Table S.7 presents an ablation study on the choice of this balance factor.
+The results indicate that a value of $0.5$ provides an optimal balance between average and worst-case identity fidelity, effectively preventing scenarios where one subject’s identity collapses while others are preserved.
+Table S.7: Ablation on Fidelity Reward $\gamma$.
+We study the effect of $\gamma$ in $\mathcal{R}_{fid}$, which balances average and minimum identity similarity.
+A value of 0.5 provides the best trade-off.
+$\gamma$ Avg FaceSim $\uparrow$ Min FaceSim $\uparrow$ Total Score $\uparrow$
+0.0 (Average only) 67.2% 45.1% 54.8%
+0.25 66.8% 50.3% 55.0%
+0.5 66.1% 53.5% 55.2%
+0.75 65.2% 52.1% 54.9%
+1.0 (Minimum only) 63.9% 51.5% 54.5%
+Figure S.7: Performance Curves of Online GRPO.
+We plot the moving average of key evaluation metrics over 800 training steps.
+Our method demonstrates simultaneous improvement in identity fidelity (FaceSim), perceptual quality (Aesthetics), and text-prompt alignment (e.g., Q-Align [wu2023qalign]), validating the effectiveness of our multi-faceted reward strategy.
+Appendix B Dataset Curation Details
+B.1 Dataset Composition and Statistics
+Our dataset is meticulously curated from three heterogeneous primary sources to ensure comprehensive coverage and high fidelity.
+The foundational component is the large-scale OpenS2V-Nexus dataset [yuan2025opens2v], which offers a wide array of authentic scenes and actions.
+To ensure the temporal coherence of the clips, we incorporate a video shot detection stage [pyscenedetect], which contributes a substantial volume of 218,230 videos, accompanied by 535,259 masked subject images and 3,098 generated counterparts.
+To further elevate the dataset’s quality, we incorporate a high-fidelity collection of professionally shot videos with detailed annotations, comprising 9,374 pristine videos and an additional 3,000 generated subject images.
+This synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, systematically increases the diversity of subject-background compositions.
 Finally, to provide rich priors for appearance synthesis, we integrate a dedicated Subject Image dataset.
 This static collection contains 14,976 images, which are further categorized into 2,877 human-centric images and 3,458 isolated clothing items.
-Figure 7: Statistics of the constructed dataset. The dataset is organized into four primary scenarios: Human, Objects, Environment, and Nature, each containing a variety of subcategories.
-Figure 8: Statistics of the constructed dataset.
+The overall composition is summarized in Table S.8, and the entire data curation pipeline is illustrated in Figure S.8.
+Table S.8: Detailed Composition of the Curated Training Dataset.
+This table provides a breakdown of the three primary data sources used to train ID-Crafter, detailing the content type, quantity, and its strategic role in the model’s training.
+Dataset Source Content Type Quantity Primary Purpose
+[…]
+[…]
+Generated Subjects 3,098 Initial subject synthesis capability
+High-Quality Internal Collection High-Fidelity Videos 9,374 Fine-tuning for video quality
+Generated Subjects 3,000 Enhancing appearance diversity
+[…]
+- Human-centric Images 2,877
+- Isolated Clothing Items 3,458
+Figure S.8: Overview of the Data Curation Pipeline.
+Our comprehensive data pipeline processes videos and images from diverse sources.
+It includes subject segmentation, high-quality reference image synthesis via an in-context learning prompt with advanced generative models, and rigorous filtering.
+This process yields a large-scale, high-fidelity dataset tailored for training identity-preserving video generation models.
+B.2 Reference Image Construction Strategies
+We analyze various strategies for constructing reference datasets in Figure S.11.
+Naive data augmentation offers limited visual diversity and is prone to generating occluded subjects, which compromises generalization to complex scenes.
+Alternatively, employing on-the-shelf image-consistent generators, such as Flux-Kontext [labs2025flux] or GPT-4o [hurst2024gpt], often introduces undesirable appearance artifacts and, most critically, fails to maintain subject fidelity.
+In contrast, our approach, which leverages a tailored in-context learning prompt for the latest Gemini-Flash-Image model (a.k.a., Nano Banana) [nanobanana], excels at subject decomposition and synthesis, yielding more faithful and varied reference images.
+Figure S.9: Statistics of the constructed dataset. The dataset is organized into four primary scenarios: Human, Objects, Environment, and Nature, each containing a variety of subcategories.
+Figure S.10: Statistics of the constructed dataset.
 (a) A word cloud illustrates the rich linguistic diversity of the reference images.
 The video duration distributions are shown for the two sources: (b) OpenS2V and (c) our curated videos.
-C.2 Comparison of Existing Datasets
-We analyze various strategies for constructing the reference datasets.
-As shown in Figure 9, Naive data augmentation offers limited visual diversity and is prone to generating occluded subjects, which compromises generalization to complex scenes.
-Alternatively, employing on-the-shelf image-consistent generators, such as Flux-Kontext (Labs et al., 2025) or even GPT-4o (Hurst et al., 2024), often introduces undesirable appearance artifacts and fails to maintain subject fidelity.
-In contrast, our approach, which leverages a tailored in-context learning prompt for the Gemini-Flash-Image model (Fortin et al., 2025), excels at subject decomposition and synthesis, yielding more faithful and varied reference images.
-Figure 9: Comparison of reference image construction strategies.
-Appendix D Details of Evaluation Metrics
+Figure S.11: Comparison of reference image construction strategies.
+In contrast to direct data augmentation or image editing models like Flux-Knotext [labs2025flux] and GPT-4O [hurst2024gpt], our data curation pipeline [Qwen2.5-VL, nanobanana] generates subject images with coherent multi-subject compositions.
+Appendix C Evaluation Metric Details
+Our evaluation protocol is principally derived from the OpenS2V-Nexus Benchmark [yuan2025opens2v].
+We adopt their prescribed prompts but substitute the proprietary models with publicly available counterparts to facilitate evaluation and ensure reproducibility.
 NexusScore: This metric quantifies the subject consistency between a generated video $\mathbf{V}=\{f_{1},\dots,f_{T}\}$ and a reference image $\mathbf{I}_{\text{ref}}$.
-It employs a two-stage pipeline that first leverages Grounded-SAM to generate a subject mask $M_{t}$ for each frame $f_{t}$.
-Subsequently, a refined CLIP-based image encoder $\mathcal{E}_{img}$ is used to compute the cosine similarity between the feature embedding of the reference image and that of the masked subject in each frame.
+It employs a two-stage pipeline: (1) Grounded-SAM [ren2024grounded] generates a subject mask $M_{t}$ for each frame $f_{t}$.
+(2) A refined CLIP-based image encoder [gal2022image] $\mathcal{E}_{img}$ computes the cosine similarity between the feature embedding of the reference image and that of the masked subject in each frame.
 This targeted approach ensures the evaluation focuses strictly on identity fidelity by mitigating background interference.
-The final score is the average similarity over all $\mathbf{T}$ frames, formulated as: $S_{\text{Nexus}}=\frac{1}{\mathbf{T}}\sum_{t=1}^{\mathbf{T}}\frac{\mathcal{E}_{\text{img}}(\mathbf{I}_{\text{ref}})\cdot\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})}{\|\mathcal{E}_{img}(\mathbf{I}_{\text{ref}})\|\|\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})\|}$ where $\odot$ denotes element-wise multiplication.
+The final score is the average similarity over all $\mathbf{T}$ frames: $S_{\text{Nexus}}=\frac{1}{\mathbf{T}}\sum_{t=1}^{\mathbf{T}}\frac{\mathcal{E}_{\text{img}}(\mathbf{I}_{\text{ref}})\cdot\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})}{\|\mathcal{E}_{img}(\mathbf{I}_{\text{ref}})\|\|\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})\|}$ where $\odot$ denotes element-wise multiplication.
 NaturalScore: This metric assesses the perceptual realism and physical plausibility of a generated video $\mathbf{V}$.
-We utilize a state-of-the-art Multimodal Large Language Model (MLLM) to perform a deep semantic analysis.
+We employ a state-of-the-art Multimodal Large Language Model (MLLM) to perform a deep semantic analysis.
 The model is presented with the video and a carefully designed prompt which instructs it to evaluate spatio-temporal consistency, identify violations of physical laws, and detect visual artifacts characteristic of generative models.
@@
 GmeScore: This metric evaluates the semantic alignment between a generated video $\mathbf{V}$ and its corresponding text prompt $\mathbf{C}$, proposed as an enhanced alternative to conventional CLIPScore.
-To overcome the limitations of prior methods in handling long-form text, GmeScore leverages the advanced text comprehension capabilities of the Qwen2-VL MLLM (Bai et al., 2025), which provides a powerful text encoder $\mathcal{E}_{text}$ and a temporally-aware video encoder $\mathcal{E}_{video}$.
+To overcome the limitations of prior methods in handling long-form text, GmeScore leverages the advanced text comprehension capabilities of the Qwen2.5-VL [Qwen2.5-VL], which provides a powerful text encoder $\mathcal{E}_{text}$ and a temporally-aware video encoder $\mathcal{E}_{video}$.
 Unlike methods that average per-frame similarities, GmeScore computes a global feature representation for the entire video, thereby capturing its holistic temporal dynamics.
-The final score is the cosine similarity between the global video embedding and the text embedding: $S_{\text{Gme}}=\frac{\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\cdot\mathcal{E}_{\text{text}}(C)}{\|\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\|\|\mathcal{E}_{\text{text}}(C)\|}$ This formulation provides a more accurate and comprehensive assessment of text-to-video relevance, especially for complex narratives.
-Appendix E More Comparison Results
-Figure 10: Prompt for MLLM-based multi-subject images generation.
-We have provided video results on the in-the-wild videos in the supplementary material. Kindly refer to our website by opening the index.html file into a modern browser.
+The final score is the cosine similarity: $S_{\text{Gme}}=\frac{\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\cdot\mathcal{E}_{\text{text}}(C)}{\|\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\|\|\mathcal{E}_{\text{text}}(C)\|}$ This formulation provides a more accurate and comprehensive assessment of text-to-video relevance, especially for complex narratives.
+Appendix D Additional Results and Analysis
+Figure S.12: Prompt for MLLM-based multi-subject images generation.
+Video results for in-the-wild inputs are available in the supplementary material Video.
+D.1 Evaluation on Multi-Subject Scenarios
+D.2 Human Preference Study
+To address concerns of reward hacking and the potential circularity of automated metrics, we conducted an extensive human preference study.
+For this evaluation, a total of 200 questionnaires were gathered from 30 participants, who compared videos generated by our model against four leading competitors across four key criteria.
+The results, summarized in Figure S.13, show a strong preference for our model in Aesthetic Quality (54%), Identity Consistency (60%), Video Quality (43%), and Motion Naturalness (65%), confirming its superior performance in generating high-quality, identity-consistent videos.
+Figure S.13: Human Preference Study Results.
+Participants were asked to choose the best video among five models (Ours and four competitors) based on four criteria.
+The chart shows the percentage of times our model was preferred.
+Our method significantly outperforms others in all categories, especially in Identity Consistency and Motion Naturalness.
+Appendix E Additional Discussion
+E.1 Computational Overhead
+The introduction of the Hierarchical Identity-Preserving Attention module adds a marginal computational overhead compared to the baseline DiT architecture.
+The primary increase stems from the three additional attention operations per layer for the initial 20 layers.
+However, since these operations are performed on the relatively small subject token sets, the impact on overall training and inference time is minimal, representing a favorable trade-off for the significant gains in identity preservation.
+The GRPO post-training phase requires additional computational resources, but it is a one-time cost that substantially enhances model performance.
+E.2 Ethics Statement
+Our work focuses on advancing controllable video generation, a technology with significant creative potential.
+We acknowledge the potential for misuse, such as the creation of misleading or harmful content.
+To mitigate these risks, we have focused our dataset curation on non-realistic, generic subjects and actions, avoiding the use of real-world public figures.
+The generated videos often contain subtle artifacts that distinguish them from real footage.
+We advocate for the development of robust detection methods for synthetic media and support the establishment of ethical guidelines for the use of generative models.
+Our model and dataset will be released to the research community to facilitate further research in this area, and we encourage responsible use.